### v1

# softMathematical software

abstract conditional independence models
Tobias Boege Department of Mathematics, KTH Royal Institute of Technology, Stockholm, Sweden post@taboege.de Janneke H. Bolt Department of Information and Computing Sciences, Utrecht University, The Netherlands & Faculty of Science, Open University of the Netherlands, Heerlen, The Netherlands j.h.bolt@uu.nl  and  Milan Studený Institute of Information Theory and Automation of the CAS, Prague, Czech Republic studeny@utia.cas.cz
(Date: February 21, 2024)
Abstract.
We introduce an algebraic concept of the frame for abstract conditional independence (CI) models, together with basic operations with respect to which such a frame should be closed: copying and marginalization. Three standard examples of such frames are (discrete) probabilistic CI structures, semi-graphoids and structural semi-graphoids. We concentrate on those frames which are closed under the operation of set-theoretical intersection because, for these, the respective families of CI models are lattices. This allows one to apply the results from lattice theory and formal concept analysis to describe such families in terms of implications among CI statements.
The central concept of this paper is that of self-adhesivity defined in algebraic terms, which is a combinatorial reflection of the self-adhesivity concept studied earlier in context of polymatroids and information theory. The generalization also leads to a self-adhesivity operator defined on the hyper-level of CI frames. We answer some of the questions related to this approach and raise other open questions.
The core of the paper is in computations. The combinatorial approach to computation might overcome some memory and space limitation of software packages based on polyhedral geometry, in particular, if SAT solvers are utilized. We characterize some basic CI families over 4 variables in terms of canonical implications among CI statements. We apply our method in information-theoretical context to the task of entropic region demarcation over 5 variables.
Key words and phrases:
self-adhesivity, conditional independence, semi-graphoid, lattice (order theory), pseudo-closed element, boolean satisfiability
2020 Mathematics Subject Classification:
62B10 (primary); 06A15, 68T27, 68V05, 05B35 (secondary)
1. Introduction
This paper is devoted to the concept of self-adhesivity in the context of abstract conditional independence (CI) structures, which generalize/approximate classic probabilistic CI structures [29]. The concepts of adhesivity and self-adhesivity have been introduced by Matúš [23] in the context of polymatroids, which are special set functions playing a crucial role in combinatorial optimization [11]. Matúš’s motivation came from information theory [35], more specifically, from the area of information-theoretical inequalities. The self-adhesivity concept for polymatroids is a kind of abstraction of the method used to derive the first non-Shannon information inequality [37], which method later became known in information-theoretical community as the so-called copy lemma [10]. The method leads to the application of tools of polyhedral geometry (and linear programming solvers) to derive new inequalities [8].
In this paper we introduce the concept of self-adhesivity for abstract CI structures, which is a combinatorial reflection of the self-adhesivity concept for polymatroids. The idea comes from [2] and has already been applied in the context of Gaussian CI structures [3]. Nonetheless, our paper puts even more emphasis on the abstract view on (the families of) CI models and utilizes the results of lattice theory [1] for this purpose. More specifically, every (finite) lattice can alternatively be described in terms of a closure operation, or in terms of (abstract) functional dependence relation [19]. Thus, one can utilize the results from the theory of formal concept analysis [13]. This leads to an analogy with mathematical logic and to the effort to describe families of CI models in an “axiomatic way”, using implications among CI statements.
We introduce the notion of an abstract frame for CI structures and give several specific examples of such CI frames. We also introduce relevant algebraic operations with respect to which these abstract CI frames should be closed: these are copying and marginalization. From the point of view of our paper, the substantial algebraic operation is the set-theoretical intersection, which ensures that, for a fixed (finite) set of variables, the family of respective CI models forms a finite lattice. Then we introduce the concept of self-adhesivity, which is relative to a CI frame. The self-adhesivity concept induces special self-adhesivity operator of the hyper-level of CI frames. This operator assigns a new CI frame to a starting CI frame and this new frame is a shrinkage of the original frame. In case the starting CI frame is an outer approximation of the classic probabilistic CI frame, the application of the self-adhesivity operator yields an even tighter outer approximation of the probabilistic CI frame.
This leads to a method to derive stronger valid probabilistic implications among CI statements on basis of previously known valid implications. The point of the story is that newly derived implications over $n$ variables are obtained on basis of known implications over a higher number of variables than $n$ through self-adhesivity.
The core of the paper are computations. Their goal was the description of several lattices of abstract CI models over a low number of variables. There are two methodologically different options: the lattices can
  • either be described by listing meet-irreducible CI models,
  • or in terms of implications among CI statements.
Our approach also allows one to derive such CI implications directly, by means of performing the computations on the combinatorial level of CI frames. To contrast with Matúš’s original concept of self-adhesivity for polymatroids, note that the dimension of the polymatroidal cone grows exponentially in $n$ and the number of CI statements grows even a little bit faster. Still, our computational methods based on SAT solvers can deal with this increase much better than the linear programming solvers could which are used in the polymatroid setup.
In the paper, a specific example of the application of our combinatorial method is given. We applied the method to the question of identification of those extreme rays of the polymatroidal cone over $5$ variables which are not entropic (= do not have probabilistic representation). It appears that most of 1319 permutational types of these extreme rays lie outside the entropic region. More specifically, it follows from our computations that at most 154 of types of such extreme rays may be entropic. The proof is entirely computational and takes only 9 minutes to execute on an average laptop. Throughout the paper we also formulate remaining open mathematical problems related to our approach.
The structure of the paper is as follows. In Section 2, we define our concept of an abstract CI frame together with corresponding algebraic operations applied to CI models. We also give three standard examples of such CI frames and recall those particular concepts from lattice theory on which our presentation is based. The concept of self-adhesion relative to such an abstract CI frame is then introduced in Section 3, together with related basic observations and two other relevant algebraic operations applied to CI models. Section 4 is then devoted to the self-adhesivity operator and the idea of its iterating. The results of our computational experiments are presented in Section 5 and their application in information-theoretical context is explained in Section 6. Concluding remarks are presented in Section 7. More detailed information about our computations, source code and resulting data files can be found on the website:
https://cinet.link/data/selfadhe-lattices
2. Basic concepts
In this section we introduce both our notational conventions and terminology. Recall that CI is an abbreviation for conditional independence.
Given a set $X$, the symbol ${\mathcal{P}}(X):=\{S:S\subseteq X\}$ will denote its power set, that is, the set of all subsets of $X$. The upper case letters like $K,L,M,N,O$ will denote finite sets of variables of interest. In our intended probabilistic interpretation, their elements correspond to random variables. Additionally, two special notational conventions will be used. Given a variable $i\in N$, the symbol $i$ will also denote the singleton set $\{i\}$ and, given two variable sets $K$ and $L$, the juxtaposition $KL$ will be a shortened notation for their union $K\cup L$. Given a bijection $\varphi:N\to M$ and $L\subseteq N$ the restriction of $\varphi$ to $L$ will be denoted by $\varphi_{|L}$, the identical bijection of $L$ onto itself by $\mbox{\rm id}_{L}$.
2.1. Abstract families of CI models
A standard elementary CI statement over $N$ is the statement claiming that, for some $K\subseteq N$ and distinct $i,j\in N\setminus K$, the variable $i$ is conditionally independent of the variable $j$ given $K$, traditionally denoted by $i\perp\!\!\!\perp j\,|\,K$. A converse conditional dependence statement (= negation) is the claim that $i$ and $j$ are conditionally dependent given $K$, traditionally denoted by $i\not\perp\!\!\!\perp j\,|\,K$. Thus, a triplet $\langle i,j|K\rangle$ of pairwise disjoint subsets of $N$ can be interpreted either as an independence statement or as a dependence statement.
Because, in this paper, we do not deal with “mixed” CI structures, involving both dependence and independence statements, we accept a convention that our sketchy notation $ij|K$ for a triplet $\langle i,j|K\rangle$ will be interpreted as the CI statement $i\perp\!\!\!\perp j\,|\,K$. Moreover, in our paper we consider only symmetric CI concepts in the sense that $i\perp\!\!\!\perp j\,|\,K$ is equivalent to $j\perp\!\!\!\perp i\,|\,K$. Every such a “symmetric” CI statement can mathematically be described by an ordered pair $\,L\,|\,K\,$ of sets $L,K\subseteq N$, where $L\cap K=\emptyset$ and $|L|=2$; specifically we have $L=\{i,j\}$. So, our second convention is that $ij|K$ means the same as $ji|K$.
Definition 2.1 (abstract CI statements, models, families, frames).
Given a finite variable set $N$, the symbol ${\mathcal{K}}(N)$ will denote the collection of all ordered pairs $\,L\,|\,K\,$, where $L,K\subseteq N$, $L\cap K=\emptyset$, and $|L|=2$.
  • The elements of ${\mathcal{K}}(N)$ will be interpreted as (CI) statements over $N$. They will usually be denoted by explicit writing: $ij|K\in{\mathcal{K}}(N)$.
  • A subset of ${\mathcal{K}}(N)$ will be interpreted as a (CI) model over $N$. Models will be denoted using the calligraphic font: ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ or ${\mathcal{M}}\in{\mathcal{P}}({\mathcal{K}}(N))$.
  • A class of (CI) models over $N$ involving the (largest possible) model ${\mathcal{K}}(N)$ will be a (CI) family over $N$. Thus, a family is a subset of the power set of ${\mathcal{K}}(N)$. Families will be denoted using the fractal font and will have the variable sets as their arguments: ${\mathcal{K}}(N)\in\mathfrak{f}(N)\subseteq{\mathcal{P}}({\mathcal{K}}(N))$.
By a (CI) frame we will understand a mapping which assigns a family over $N$ to every finite (variable) set $N$. Frames will be denoted using the fractal font without an argument: mapping $\mathfrak{f}:\,N\mapsto\mathfrak{f}(N)\in{\mathcal{P}}({\mathcal{P}}({\mathcal{K}}(N)))$.
Given two CI frames $\mathfrak{f}_{1}$ and $\mathfrak{f}_{2}$, the record $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$ means that $\mathfrak{f}_{1}(N)\subseteq\mathfrak{f}_{2}(N)$ for any (finite) variable set $N$.
The reader can recognize a clear hierarchy: a (CI) statement is an element of ${\mathcal{K}}(N)$, a model is an element of the power set of ${\mathcal{K}}(N)$ and a family is an element of the power set of the power set of ${\mathcal{K}}(N)$ (= the double power set). [^1] A frame corresponds to what was called a “property” in the PhD thesis [2, § 2.2].
Note that ${\mathcal{K}}(N)=\emptyset$ if $N$ is empty or a singleton. In particular, only one (CI) model over $N$ exists if $|N|\leq 1$, namely the empty set/model. Moreover, only one (CI) family over $N$ exists if $|N|\leq 1$ because of our explicit requirement that ${\mathcal{K}}(N)\in\mathfrak{f}(N)$ in the definition of a family.
The next step is to introduce some “algebraic” operations with abstract CI models. The goal is to study various frames which are closed under such operations. Later, in Section 2.2, we give specific examples of abstract CI frames we have in mind.
Definition 2.2 (copying, marginalization, intersection, duality).
  • Any bijection $\varphi:N\to M$ between two variable sets $N$ and $M$ can naturally be extended to a bijection between their power sets and to a bijection between ${\mathcal{K}}(N)$ and ${\mathcal{K}}(M)$: put $\varphi(ij|K)\,:=\,\varphi(ij)|\varphi(K)$ for any $ij|K\in{\mathcal{K}}(N)$.
  • Any model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ is then transformed by $\varphi$ to a model over $M$: put $\varphi({\mathcal{M}}):=\{\,\varphi(ij|K)\,:\ ij|K\in{\mathcal{M}}\,\}\subseteq{\mathcal{K}}(M)$. This is the $\varphi$-copy of ${\mathcal{M}}$. We say that a CI frame $\mathfrak{f}$ is closed under copying if, for any model ${\mathcal{M}}\in\mathfrak{f}(N)$ and any bijection $\varphi:N\to M$, one has $\varphi({\mathcal{M}})\in\mathfrak{f}(M)$.
  • Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$ and  $M\subseteq N$, its marginal for $M$, denoted by ${\mathcal{M}}^{\downarrow M}$, is simply its restriction to ${\mathcal{K}}(M)$: put ${\mathcal{M}}^{\downarrow M}:={\mathcal{M}}\cap{\mathcal{K}}(M)$. A CI frame $\mathfrak{f}$ is closed under marginalization if, for any model ${\mathcal{M}}\in\mathfrak{f}(N)$ and any subset $M\subseteq N$, one has ${\mathcal{M}}^{\downarrow M}\in\mathfrak{f}(M)$.
  • Given two models ${\mathcal{M}},{\mathcal{R}}\subseteq{\mathcal{K}}(N)$ over the same set $N$, their (set-theoretical) intersection is the model ${\mathcal{M}}\cap{\mathcal{R}}$, again a model over $N$. A CI frame $\mathfrak{f}$ will be called closed under intersection if, for any two models ${\mathcal{M}},{\mathcal{R}}\in\mathfrak{f}(N)$, one has ${\mathcal{M}}\cap{\mathcal{R}}\in\mathfrak{f}(N)$.
  • Given a variable set $N$, a duality mapping from ${\mathcal{K}}(N)$ to ${\mathcal{K}}(N)$ is as follows: $ij|K\in{\mathcal{K}}(N)$ is assigned the pair $(ij|K)^{\star}:=ij|L$, where $L\,:=\,N\setminus(K\cup ij)$. Given a model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$, the dual model ${\mathcal{M}}^{\star}$ is its image by this mapping. A CI frame $\mathfrak{f}$ is closed under duality if ${\mathcal{M}}\in\mathfrak{f}(N)$ implies ${\mathcal{M}}^{\star}\in\mathfrak{f}(N)$.
Since ${\mathcal{K}}(N)=\emptyset$ if $|N|\leq 1$, the concepts of copying and duality make real sense only in case $|N|\geq 2$. On the other hand, to introduce later concepts from Section 3 one needs to define marginalization to sets $M$ with $|M|\leq 1$.
Two basic “algebraic” operations with abstract CI models are copying and marginalization. All frames we consider in this paper are closed under these two operations. Most of the considered CI frames are also closed under intersection, which implies that the respective families of models are lattices (see Section 2.3). A more complicated adhesivity concept will be discussed later (Section 3).
2.2. Examples of abstract CI frames
There are several distinct frames we consider in this paper.
2.2.1. Discrete probabilistic CI models
A prototypical example of a CI frame is the frame of discrete probabilistic CI structures. Here we define them specifically.
[…]
$\forall\,x\in\mbox{\sf X}_{N}\qquad p_{ijK}(x)\cdot p_{K}(x)=p_{iK}(x)\cdot p_{jK}(x)\,.$
The CI model induced by $\xi$ is the set ${\mathcal{M}}_{\mbox{\scriptsize\boldmath$\xi$}}:=\{\,ij|K\in{\mathcal{K}}(N)\,:\ i\perp\!\!\!\perp j\,|\,K\,\,[\mbox{\boldmath$\xi$}]\,\}$. The family of discrete probabilistic models over $N$ is then
$\mathfrak{pr}\,(N)\,:=\,\{\,{\mathcal{M}}_{\mbox{\scriptsize\boldmath$\xi$}}\,:\ \mbox{\boldmath$\xi$}~{}~{}\mbox{\rm is a discrete random vector over $N$}\,\}\,.$
This defines the discrete probabilistic CI frame $\mathfrak{pr}$.
Lemma 2.3.
$\mathfrak{pr}$ is closed under copying, marginalization, and intersection.
Proof 2.4.
[…]
On the other hand, the CI frame $\mathfrak{pr}$ is not closed under duality. Recall a counter-example from [21, end of § 5].
Example 2.5.
Given $N:=\{a,b,c,d\}$ and :=${\mathcal{M}}:=\{\,ab|cd,\,ab|d,\,ab|c,\,cd|\emptyset\,\}$, one has ${\mathcal{M}}\in\mathfrak{pr}\,(N)$ and =${\mathcal{M}}^{\star}=\{\,ab|\emptyset,\,ab|c,\,ab|d,\,cd|ab\,\}\not\in\mathfrak{pr}\,(N)$. Indeed, a binary random vector uniformly distributed on four special configurations
$(0,0,0,0)\quad(0,1,0,1)\quad(0,1,1,0)\quad(1,1,1,1)$
of values for $(x_{a},x_{b},x_{c},x_{d})$ implies the former fact. The latter fact follows from the property (I:13) in [31, § V.B] saying that the implication
$[\,\,a\perp\!\!\!\perp b\,|\,\emptyset~{}~{}\&~{}~{}a\perp\!\!\!\perp b\,|\,c~{}~{}\&~{}~{}a\perp\!\!\!\perp b\,|\,d~{}~{}\&~{}~{}c\perp\!\!\!\perp d\,|\,ab\,\,]\quad\Rightarrow\quad a\perp\!\!\!\perp b\,|\,cd$
holds in the context of discrete probabilistic CI structures.
2.2.2. Semi-graphoids
The concept of a semi-graphoid was introduced by Pearl [26] with the aim to formalize fundamental properties of probabilistic CI pinpointed earlier by Dawid [9]. Here we give an equivalent definition from [20, § 5] adapted to our situation.
A semi-graphoid over a variable set $N$ is a subset ${\mathcal{M}}$ of ${\mathcal{K}}(N)$ satisfying
$ij|K,\,i\ell|jK\in{\mathcal{M}}~{}\Leftrightarrow~{}i\ell|K,\,ij|\ell K\in{\mathcal{M}}\,.$
Then we put
$\mathfrak{sg}\,(N)\,:=\,\{\,{\mathcal{M}}\subseteq{\mathcal{K}}(N)\,:\ \mbox{\rm${\mathcal{M}}$ is a semi-graphoid over $N$}\,\}\,,$
which defines the semi-graphoidal CI frame $\mathfrak{sg}$.
Lemma 2.6.
The frame $\mathfrak{sg}$ is closed under copying, marginalization, intersection, and duality.
Proof 2.7.
This basically follows from the above definition of a semi-graphoid.
Well-known facts are that $\mathfrak{pr}\subseteq\mathfrak{sg}$ and $\mathfrak{sg}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$. The simplest example that $\mathfrak{sg}(N)\neq\mathfrak{pr}(N)$ in case $|N|\geq 4$ is as follows.
Example 2.8.
Given $N:=\{a,b,c,d\}$ and :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b\,\}\in\mathfrak{sg}(N)$, one has ${\mathcal{M}}\not\in\mathfrak{pr}(N)$ by property (E:1) in [31, § V.A] saying that the equivalence
$[\,\,a\perp\!\!\!\perp b\,|\,c~{}~{}\&~{}~{}a\perp\!\!\!\perp c\,|\,d~{}~{}\&~{}~{}a\perp\!\!\!\perp d\,|\,b\,\,]~{}\Leftrightarrow~{}[\,\,a\perp\!\!\!\perp b\,|\,d~{}~{}\&~{}~{}a\perp\!\!\!\perp c\,|\,b~{}~{}\&~{}~{}a\perp\!\!\!\perp d\,|\,c\,\,]$
holds in the context of discrete probabilistic CI structures.
Definition 2.9 (global CI statement).
Let $N$ be a variable set and $I,J,K\subseteq N$ pairwise disjoint subsets of it. Given a CI model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over $N$, we will write $I\perp\!\!\!\perp J\,|\,K\,\,[{\mathcal{M}}]$ if
(1) $\forall\,i\in I\quad\forall\,j\in J\quad\forall\,L:K\subseteq L\subseteq IJK\setminus ij\qquad ij|L\in{\mathcal{M}}\,.$
We interpret this as a global CI statement encoded within the model ${\mathcal{M}}$. In our examples and catalogues we also use a shorthand $[I,J|K]$ to denote the list of (elementary) CI statements described in (1).
Despite the global CI statement is formally defined relative to any CI model ${\mathcal{M}}$, the concept itself makes sense only when ${\mathcal{M}}$ is a semi-graphoid. This relates to two alternative modes for describing semi-graphoids: in terms of either global or local (= elementary) CI statements; see [20, § 5] or [29, Lemma  2.2].
Remark. This is to elucidate possible alternative ways to formalize CI models. An equivalent way to the one from Definition 2.1 is to replace ${\mathcal{K}}(N)$ by the collection ${\mathcal{T}}(N)$ of all triplets $\langle I,J|K\rangle$ of pairwise disjoint subsets $I,J,K$ of $N$ and define (global) CI models as subsets of ${\mathcal{T}}(N)$. These two equivalent approaches are then related through (the concept introduced in) Definition 2.9.
A non-equivalent approach is to replace ${\mathcal{K}}(N)$ by the collection ${\mathcal{K}}^{\bullet}(N)$ of ordered pairs $\,L\,|\,K\,$, where $L,K\subseteq N$, $L\cap K=\emptyset$, and $1\leq|L|\leq 2$. Any pair $i\,|\,K$ then corresponds to a functional dependence statement claiming that $i\in N\setminus K$ is “determined” by $K\subseteq N$, interpreted as a specific CI statement $i\perp\!\!\!\perp i\,|\,K$ (= $i$ is conditionally independent of itself given $K$). The subsets of ${\mathcal{K}}^{\bullet}(N)$ are then augmented CI models over $N$ [31, § II.B] involving functional dependencies, represented in the local/elementary mode. This wider approach can equivalently be described by replacing ${\mathcal{K}}^{\bullet}(N)$ with the collection ${\mathcal{T}}^{\bullet}(N)$ of all triplets $\langle I,J|K\rangle$ of subsets $I,J,K$ of $N$. The subsets of ${\mathcal{T}}^{\bullet}(N)$ can be viewed as augmented CI models, represented in the global mode. However, we do not consider augmented CI models with functional dependence statements in this paper.
2.2.3. Structural semi-graphoids
The concept of a structural semi-graphoid was introduced in [28] as a semi-graphoid induced in a particular way by certain subsets of ${\mathbb{Z}}^{{\mathcal{P}}(N)}$. The original complicated definition was shown in [29, Corollary 5.3] to be equivalent to a much simpler condition that underlies our definition below.
Definition 2.10 (supermodular function).
Given a set function $m:{\mathcal{P}}(N)\to{\mathbb{R}}$ and arbitrary subsets $I,J,K\subseteq N$, we are going to use the following special notation for a particular value:
$\Delta m\,\langle I,J|K\rangle~{}:=~{}m(IJK)+m(K)-m(IK)-m(JK)\,.$
The supermodular difference expression for $ij|K\in{\mathcal{K}}(N)$ is a special case of it:
$\Delta m(ij|K):=\Delta m\,\langle i,j|K\rangle=m(ijK)+m(K)-m(iK)-m(jK)\,.$
The function $m$ is called supermodular if $\Delta m(ij|K)\geq 0$ for any $ij|K\in{\mathcal{K}}(N)$, which requirement is equivalent to $\Delta m\,\langle I,J|K\rangle\geq 0$ for pairwise disjoint subsets $I,J,K\subseteq N$, or to the well-known standard supermodularity condition
$m(A\cup B)+m(A\cap B)\geq m(A)+m(B)\qquad\mbox{for any $A,B\subseteq N$.}$
The model ${\mathcal{M}}^{m}:=\{\,ij|K\in{\mathcal{K}}(N)\,:\ \Delta m(ij|K)=0\,\}$, where $m:{\mathcal{P}}(N)\to{\mathbb{R}}$ is supermodular, is the structural semi-graphoid induced by $m$. The class of structural semi-graphoids over $N$ is
$\mathfrak{st}\,(N)\,:=\,\{\,{\mathcal{M}}^{m}\,:\ ~{}\mbox{\rm$m$ is a supermodular function on ${\mathcal{P}}(N)$}\,\}\,.$
This defines the structural CI frame $\mathfrak{st}$.
Lemma 2.11.
The frame $\mathfrak{st}$ is closed under copying, marginalization, intersection, and duality.
Proof 2.12.
Given a bijection $\varphi:N\to M$ and a set function $m:{\mathcal{P}}(N)\to{\mathbb{R}}$, one has $\Delta m(ij|K)=\Delta(m\circ\varphi_{-1})\,\varphi(ij|K)$ for any $ij|K\in{\mathcal{K}}(N)$, where $\varphi$ is naturally extended to the power set of $N$ and ${\mathcal{K}}(N)$. Hence, provided that $m$ is a supermodular function on ${\mathcal{P}}(N)$, the composition $m\circ\varphi_{-1}$ is a supermodular function on ${\mathcal{P}}(M)$ and the $\varphi$-copy of ${\mathcal{M}}^{m}$ is ${\mathcal{M}}^{m\circ\varphi_{-1}}$.
If $M\subseteq N$ then the restriction of a supermodular function on ${\mathcal{P}}(N)$ to ${\mathcal{P}}(M)$ is supermodular, which implies the claim about the marginalization.
Given two set functions $m^{1}:{\mathcal{P}}(N)\to{\mathbb{R}}$ and $m^{2}:{\mathcal{P}}(N)\to{\mathbb{R}}$, one has $\Delta(m^{1}+m^{2})(ij|K)=\Delta m^{1}(ij|K)+\Delta m^{2}(ij|K)$ for any $ij|K\in{\mathcal{K}}(N)$, which implies that if $m^{1}$ and $m^{2}$ are supermodular then $m:=m^{1}+m^{2}$ is supermodular. The non-negativity of values $\Delta m^{1}(ij|K)$ and $\Delta m^{2}(ij|K)$ allows one to deduce that ${\mathcal{M}}^{m}={\mathcal{M}}^{m^{1}}\cap{\mathcal{M}}^{m^{2}}$.
Consider a self-inverse mapping $\iota:{\mathcal{P}}(N)\to{\mathcal{P}}(N)$ defined by $\iota(S):=N\setminus S$ for any $S\subseteq N$. The formula $\Delta m(ij|K)=\Delta(m\circ\iota)(ij|K)^{\star}$ valid for any $m:{\mathcal{P}}(N)\to{\mathbb{R}}$ and $ij|K\in{\mathcal{K}}(N)$, implies that $m$ is supermodular if and only if $m\circ\iota$ is supermodular and that $({\mathcal{M}}^{m})^{\star}={\mathcal{M}}^{m\circ\iota}$.
Basic facts are that $\mathfrak{pr}\subseteq\mathfrak{st}\subseteq\mathfrak{sg}$ and $\mathfrak{sg}(N)=\mathfrak{st}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$; for $\mathfrak{pr}\subseteq\mathfrak{st}$ see [29, § 2.3.4]. The strict inclusion $\mathfrak{st}(N)\subset\mathfrak{sg}(N)$ in case $|N|=4$ is evidenced by the model :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b\,\}$ from Example 2.8, because the property (E:1) mentioned there holds for structural semi-graphoids.
The strict inclusion $\mathfrak{pr}(N)\subset\mathfrak{st}(N)$ in case $|N|=4$ is evidenced by the dual model :=${\mathcal{M}}^{\star}:=\{\,ab|\emptyset,\,ab|c,\,ab|d,\,cd|ab\,\}$ from Example 2.5, which is induced by the following supermodular function $m$ on $N=\{a,b,c,d\}$:
$m(S)~{}:=~{}\left\{\begin{array}[]{cl}12&\mbox{if $S=abcd$,}\\ 7&\mbox{if $S=acd$ or $S=bcd$,}\\ 6&\mbox{if $S=abc$ or $S=abd$,}\\ 3&\mbox{if $S\in\{\,ac,ad,bc,bd,cd\,\}$,}\\ 0&\mbox{otherwise,}\end{array}\right.~{}~{}\qquad\mbox{for $S\subseteq N$.}$
Remarks.  This is to remind alternative views.
  1. (1)
    A set function $h:{\mathcal{P}}(N)\to{\mathbb{R}}$ is submodular if it fulfils the converse inequalities to the supermodular ones :
    $h(A\cup B)+h(A\cap B)\leq h(A)+h(B)\qquad\mbox{for any $A,B\subseteq N$,}$
    which is another way of saying that $-h$ is supermodular. The rank function of a polymatroid over $N$ is a submodular function $h$ on ${\mathcal{P}}(N)$ which satisfies $h(\emptyset)=0$ and is non-decreasing: $h(A)\leq h(B)$ if $A\subseteq B\subseteq N$.
    One can equivalently introduce structural semi-graphoids as CI models induced by submodular functions or by polymatroids in an analogous way. Clearly, it is just a matter of taste and convention whether the supermodular view or the submodular view is chosen.
  2. (2)
    The concept of a structural semi-graphoid can be viewed as a reduced version of the concept of a semi-matroid from [24, § 2]. The difference is that semi-matroids involve also functional dependence statements, interpreted as special CI statements $i\perp\!\!\!\perp i\,|\,K$ for $i\in N$, $K\subseteq N\setminus\{i\}$. But otherwise semi-matroids are defined analogously: they are induced by polymatroids through vanishing of the respective submodular difference expressions.
    The polymatroidal attitude allows one to emphasize the links to matroid theory [25] because polymatroids generalize matroids. Specifically, a polymatroidal rank function $h$ on ${\mathcal{P}}(N)$ corresponds to a matroid over $N$ if and only if it takes values in ${\mathbb{Z}}$ and additionally satisfies $h(I)\leq|I|$ for any $I\subseteq N$.
2.3. Lattices of models and closure operations
Given a CI frame $\mathfrak{f}$ closed under intersection and a fixed variable set $N$, the poset $(\mathfrak{f}(N),\subseteq)$ appears to be a finite lattice. Such a lattice can be characterized in two methodologically different ways:
  • in terms of irreducible models (= an inner description in the sense that the characterization only depends on the lattice itself),
  • in terms of implicational generators (= an outer description defined in case the lattice is specifically embedded into a Boolean lattice).
In this section we recall relevant concepts from lattice theory [1].
Given a set $X$, a Moore family of subsets of $X$, alternatively called a closure system (for $X$), is a collection ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$ of subsets of $X$ which is closed under (arbitrary) set intersection and includes the set $X$ itself:
${\mathcal{D}}\subseteq{\mathcal{F}}~{}\Rightarrow~{}\bigcap{\mathcal{D}}\in{\mathcal{F}}\quad\mbox{with a convention that\, $\bigcap{\mathcal{D}}=X$ for ${\mathcal{D}}=\emptyset$}.$
An elementary observation is that any Moore family $({\mathcal{F}},\subseteq)$, equipped with the set inclusion relation, is a complete lattice [1, Theorem 2 in § V.1]. [^2] This is a universal construction in the sense that any lattice is isomorphic to a lattice based on a Moore family and any finite lattice is isomorphic to a finite Moore family lattice. These conclusions can be derived from observations in [1, § V.9]. In our treatment the intended setting is $X:={\mathcal{K}}(N)$ for a finite variable set $N$. Thus, we have the following observation.
Corollary 1.
If a CI frame $\mathfrak{f}$ is closed under intersection then, for each set $N$, the family $(\mathfrak{f}(N),\subseteq)$ is a finite lattice in which infimum means intersection. If $\mathfrak{f}$ is also closed under copying and $|N|=|M|$ then $(\mathfrak{f}(N),\subseteq)$ and $(\mathfrak{f}(M),\subseteq)$ are order-isomorphic and permutations of $N$ yield automorphisms of $(\mathfrak{f}(N),\subseteq)$.
Proof 2.13.
As $\mathfrak{f}(N)$ is a Moore family, this is evident from what was said before.
Let $({\mathcal{L}},\preceq)$ be a (finite) lattice with the greatest element denoted by 1 and the meet/infimum operator by $\inf$. Let $\prec$  denote the strict order relation. Given $\lambda,\tau\in{\mathcal{L}}$, one says that $\tau$ covers $\lambda$ and writes $\lambda\prec\!\!\!\cdot\,\,\,\tau$ if $\lambda\prec\tau$ and there is no $\kappa\in{\mathcal{L}}$ with $\lambda\prec\kappa\prec\tau\/$. A coatom of ${\mathcal{L}}$ is a sub-maximal element in ${\mathcal{L}}$, that is, $\tau\in{\mathcal{L}}$ satisfying $\tau\prec\!\!\!\cdot\,\,\,\mbox{\bf 1}$. An element $\lambda\in{\mathcal{L}}$ is called (meet) irreducible if $\lambda\neq\inf\,\{\,\tau\in{\mathcal{L}}\,:\ \lambda\prec\tau\,\}$. [^3] Note that 1 is not irreducible by this definition: it is the infimum of the empty set. A well-known fact is that the set ${\mathcal{I}}$ of (meet) irreducible elements in a finite lattice $({\mathcal{L}},\preceq)$ is the least set determining all elements of ${\mathcal{L}}$ by means of the infimum operation in the following sense:
$\forall\,\kappa\in{\mathcal{L}}\quad\exists\,{\mathcal{S}}\subseteq{\mathcal{I}}\qquad\mbox{such that}~{}~{}\kappa=\inf\,\{\,\tau\,:\ \tau\in{\mathcal{S}}\,\}\,.$
The coatoms are standard examples of irreducible elements in ${\mathcal{L}}$. In general, a lattice can have other irreducible elements. A lattice $({\mathcal{L}},\preceq)$ is called coatomistic if every element of ${\mathcal{L}}$ is the infimum of a set of coatoms, which is another way of saying that there is no irreducible element in ${\mathcal{L}}$ except the coatoms.
Example 2.14.
Put $X:=\{x,y,z,w\}$ and consider a Moore family
:=${\mathcal{F}}:=\{\,\,\{x,y,z,w\},\{x,y,z\},\{x,y,w\},\{x,y\},\{x\}\,\,\}\,,$
of subsets of $X$, depicted in Figure 1 in the form of its Hasse diagram. The elements $\{x,y,z\}$, $\{x,y,w\}$ and $\{x\}$ of the lattice $({\mathcal{F}},\subseteq)$ are then irreducible, but only the elements $\{x,y,z\}$ and $\{x,y,w\}$ are its coatoms. Thus, $({\mathcal{F}},\subseteq)$ is a finite lattice which is not coatomistic.
Figure 1. A Moore family $\{\,\{x,y,z,w\},\{x,y,z\},\{x,y,w\},\{x,y\},\{x\}\,\}$.
Thus, the first way to characterize a finite lattice $({\mathcal{F}},\subseteq)$, where ${\mathcal{F}}$ is a Moore family of subsets of $X$, is to give the complete list of irreducible elements in ${\mathcal{F}}$. We now apply this to the case $X:={\mathcal{K}}(N)$ for a finite variable set $N$ and ${\mathcal{F}}:=\mathfrak{f}(N)$ being a suitable family (of CI models) over $N$.
Corollary 2.
If a CI frame $\mathfrak{f}$ is closed under copying and intersection then, for each variable set $N$, the irreducible sets in $(\mathfrak{f}(N),\subseteq)$ (= irreducible models) fall into permutational equivalence classes (= relative to permutations of $N$).
Proof 2.15.
By Corollary 1, any permutation $\pi:N\to N$ defines an order automorphism of $(\mathfrak{f}(N),\subseteq)$, namely the natural extension of $\pi$ to ${\mathcal{K}}(N)$. Such a transformation preserves the irreducibility of an element of $\mathfrak{f}(N)$.
Example 2.16.
(lattice of semi-graphoids over 3 variables)
Let us take $N:=\{a,b,c\}$, $X:={\mathcal{K}}(N)$, and put ${\mathcal{F}}:=\mathfrak{sg}\,(N)$. The lattice $(\mathfrak{sg}\,(N),\subseteq)$ has 22 elements in this case and they break into 10 permutational equivalence classes. There are 5 (meet) irreducible elements in $\mathfrak{sg}\,(\{a,b,c\})$ which fall down into 3 permutational types:
  • $\{\,ab|\emptyset,\,ac|\emptyset,\,bc|\emptyset\,\}$,
  • $\{\,ab|c,\,ac|b,\,bc|a\,\}$,
  • $\{\,ab|\emptyset,\,ab|c,\,ac|\emptyset,\,ac|b\,\}$ (with 2 additional permutational copies).
All these elements are coatoms; in particular, the lattice is coatomistic.
A kind of dual description of a finite lattice is based on the following concept. A closure operation on subsets of $X$ is a mapping $\mbox{\rm cl}:{\mathcal{P}}(X)\to{\mathcal{P}}(X)$ which is
  • extensive: $Y\subseteq\mbox{\rm cl}\,(Y)$ for $Y\subseteq X$,
  • isotone: $Y\subseteq Z\subseteq X\Rightarrow\mbox{\rm cl}\,(Y)\subseteq\mbox{\rm cl}\,(Z)$, and
  • idempotent: $\mbox{\rm cl}\,(\mbox{\rm cl}\,(Y))=\mbox{\rm cl}\,(Y)$ for $Y\subseteq X$.
A set $Y\subseteq X$ is then called closed with respect to cl if $Y=\mbox{\rm cl}\,(Y)$. A basic fact is that the collection of subsets of $X$ closed with respect to cl is a Moore family and any Moore family can be defined in this way [1, Theorem 1 in § V.1]. Specifically, given a Moore family ${\mathcal{F}}$ of subsets of $X$, the formula
$\mbox{\rm cl}_{{\mathcal{F}}}(Y)\,:=\,\bigcap\,\{\,Z\subseteq X\,:\,Y\subseteq Z\in{\mathcal{F}}\,\}\quad\mbox{\rm for}~{}Y\subseteq X\,,$
defines a closure operation on subsets of $X$ having ${\mathcal{F}}$ as the collection of closed sets with respect to $\mbox{\rm cl}_{{\mathcal{F}}}$; see [13, Theorem 1 in § 0.3]. Thus, for any $Y\subseteq X$, the closure $\mbox{\rm cl}_{{\mathcal{F}}}(Y)$ of $Y$ relative to ${\mathcal{F}}$ is the smallest element of ${\mathcal{F}}$ containing $Y$.
Example 2.17.
Consider the Moore family ${\mathcal{F}}$ from Example 2.14. Then, for instance, $\mbox{\rm cl}_{{\mathcal{F}}}(\emptyset)=\{x\}=\mbox{\rm cl}_{{\mathcal{F}}}(\{x\})$, $\mbox{\rm cl}_{{\mathcal{F}}}(\{y\})=\{x,y\}$, and $\mbox{\rm cl}_{{\mathcal{F}}}(\{x,z\})=\{x,y,z\}$.
The closure operation can also be interpreted as a functional dependence relation on subsets of $X$, as explained in [19]. This is a binary relation $\to$ on ${\mathcal{P}}(X)$ satisfying three axioms (for $Y,Z,W\subseteq X$):
  • reflexivity: $Z\subseteq Y~{}\Rightarrow~{}Y\to Z$,
  • transitivity: $[\,Y\to Z~{}\&~{}Z\to W\,]~{}\Rightarrow~{}Y\to W$,
  • union: $[\,Y\to Z~{}\&~{}Y\to W\,]~{}\Rightarrow~{}Y\to Z\cup W$.
Together, these axioms are equivalent to the well-known Armstrong axioms for functional dependence relations in database theory. Given a closure operation cl on subsets of $X$, define $Y\to Z$ as $Z\subseteq\mbox{\rm cl}\,(Y)$. Given a functional dependence relation $\to$, put $\mbox{\rm cl}\,(Y):=\bigcup\,\{\,Z\subseteq X\,:\ Y\to Z\,\}$, which closure operation induces the original functional dependence relation $\to$.
Example 2.18.
Let us consider again the Moore family ${\mathcal{F}}$ from Example 2.14 and its induced closure operation $Y\mapsto\mbox{\rm cl}_{{\mathcal{F}}}(Y)$ for $Y\subseteq X$. The respective functional dependence relation is quite huge. It certainly contains 16 implications saying that subsets of $X$ imply their closures:
$\emptyset\to\{x\},~{}~{}\{x\}\to\{x\},~{}~{}\{y\}\to\{x,y\},~{}~{}\{z\}\to\{x,y,z\},~{}~{}\{w\}\to\{x,y,w\},$
[…]
Besides these, it also contains their natural consequences because of inclusion: for instance, it contains the implication $\{y,z,w\}\to\{x,w\}$ which follows from the inclusion $\{x,w\}\subseteq\mbox{\rm cl}_{{\mathcal{F}}}(\{y,z,w\})$ (use reflexivity and transitivity for this).
The second way to characterize a finite lattice $({\mathcal{F}},\subseteq)$, where ${\mathcal{F}}$ is a Moore family of subsets of $X$, is to give a suitable generator of the respective functional dependence relation $\to$. Before explaining the details we wish the reader to realize/observe that there is a straightforward relationship (= correspondence) between the functional dependence relations $\to$ and Moore families ${\mathcal{F}}$ [^4] A one-to-one assignment between a Moore family and a functional dependence relation has already been established through their shared closure operation. We now re-interpret this as a kind of duality relationship and avoid the concept of a closure operation.
Specifically, any element $(Y,Z)$ of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ can be interpreted as a formal implication $Y\to Z$, meaning that the set $Y$ of antecedents implies (all elements of) the set $Z$ of consequents. One can introduce a binary relation between elements of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and ${\mathcal{P}}(X)$ as follows:
$\forall\,(Y\to Z)\in{\mathcal{P}}(X)\times{\mathcal{P}}(X)~{}~{}\forall\,F\in{\mathcal{P}}(X)$
$(Y\to Z)\odot F\quad:=\quad[\,\,Y\setminus F\neq\emptyset~{}~{}\mbox{or}~{}~{}Z\subseteq F\,\,],$
that is, $[\,Y\subseteq F~{}\Rightarrow~{}Z\subseteq F\,]$, meaning that $Y\to Z$ is valid for $F$.
The next step is to define the so-called Galois connections between subsets of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and subsets of ${\mathcal{P}}(X)$:
$\!\!{\mathbb{G}}\subseteq{\mathcal{P}}(X)\times{\mathcal{P}}(X)\,\mapsto\,{\mathbb{G}}^{\triangleright}:=\{\,F\in{\mathcal{P}}(X):\,\forall\,(Y\to Z)\in{\mathbb{G}}~{}\,(Y\to Z)\odot F\,\},$
${\mathcal{F}}\subseteq{\mathcal{P}}(X)\,\mapsto\,{{\mathcal{F}}}^{\triangleleft}:=\{\,(Y\to Z)\in{\mathcal{P}}(X)\times{\mathcal{P}}(X):\,\forall\,F\in{\mathcal{F}}~{}\,(Y\to Z)\odot F\,\}\,.$
Well-known classic results in lattice theory on Galois connections [1, § V.7], see also specifically [19, Theorem], say the following:
  • the mapping ${\mathbb{G}}\mapsto{\mathbb{G}}^{\triangleright\triangleleft}$ is a closure operation on subsets of (a much larger set) ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and ${\mathbb{G}}^{\triangleright\triangleleft}$ is then the smallest functional dependence relation (on subsets of $X$) containing ${\mathbb{G}}$,
  • the mapping ${\mathcal{F}}\mapsto{{\mathcal{F}}}^{\triangleleft\triangleright}$ is a closure operation on subsets of ${\mathcal{P}}(X)$ and ${{\mathcal{F}}}^{\triangleleft\triangleright}$ is then the smallest Moore family (of subsets of $X$) containing ${\mathcal{F}}$,
  • a system ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$ is a Moore family if and only if $[\,\exists\,{\mathbb{G}}\subseteq{\mathcal{P}}(X)\times{\mathcal{P}}(X)\,:\ {\mathcal{F}}={\mathbb{G}}^{\triangleright}\,]$.
We say that ${\mathbb{G}}\subseteq{\mathcal{P}}(X)\times{\mathcal{P}}(X)$ is an implicational generator of a Moore family ${\mathcal{F}}$ if ${\mathcal{F}}={\mathbb{G}}^{\triangleright}$. Elementary facts on Galois connections allow one to observe that an equivalent condition is ${{\mathcal{F}}}^{\triangleleft}={\mathbb{G}}^{\triangleright\triangleleft}$. Hence, ${\mathbb{G}}$ is an implicational generator of a Moore family ${\mathcal{F}}$ if and only if both ${\mathbb{G}}\subseteq{{\mathcal{F}}}^{\triangleleft}$ (soundness) and $|{\mathbb{G}}^{\triangleright}|=|{{\mathcal{F}}}|$ (completeness).
Example 2.19.
Consider $X:=\{x,y,z,w\}$ and put
${\mathbb{G}}:=\{~{}\emptyset\to\{x\},~{}\{z\}\to\{y\},~{}\{x,w\}\to\{x,y,w\}~{}\}\,.$
Then ${\mathbb{G}}^{\triangleright}$ equals to the family ${\mathcal{F}}$ from Example 2.14. Hence, ${\mathbb{G}}$ is one of several possible implicational generators of ${\mathcal{F}}$, a minimal one in sense of inclusion.
To introduce a prominent implicational generator one needs specific concepts from (theory of) formal concept analysis; see [12, Chapter  3] or [34, § 1.3.2]. [^5] In general, there is no unique smallest implicational generator is sense of inclusion.
Definition 2.20 (pseudo-closed sets, canonical basis).
We say that $Q\subseteq X$ is pseudo-closed relative to a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$ if
  • $Q\not\in{\mathcal{F}}$ (= $Q$ is not closed),
  • $\forall\,R\subset Q~{}\mbox{\rm pseudo-closed}~{}~{}\mbox{\rm cl}_{{\mathcal{F}}}(R)\subseteq Q$ (then $\mbox{\rm cl}_{{\mathcal{F}}}(R)\subset Q$ because $Q\not\in{\mathcal{F}}$).
The class of pseudo-closed subsets relative to ${\mathcal{F}}$ will be denoted by $\wp({\mathcal{F}})$. The canonical implicational basis for a Moore family ${\mathcal{F}}$ is as follows:
${\mathbb{G}}_{{\mathcal{F}}}~{}:=~{}\{\,(\,Q\to\mbox{\rm cl}_{{\mathcal{F}}}(Q)\setminus Q\,)\,:\ Q\in\wp({\mathcal{F}})\,\}\,\subseteq\,{\mathcal{P}}(X)\times{\mathcal{P}}(X)\,.$
The recursive definition of pseudo-closedness of a set $Q$ is well-formed as it depends only on the previously defined pseudo-closedness of strict subsets of $Q$. The canonical implicational basis, alternatively named Duquenne–Guigues basis, was introduced already in the 1980s [15]. We have also implemented an algorithm to compute the canonical basis ${\mathbb{G}}_{\mathcal{F}}$ for a given Moore family ${\mathcal{F}}$; its idea is described in Appendix A.
Every inclusion-minimal non-closed subset, that is, $Q\in({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$, is always pseudo-closed since all proper subsets of $Q$ are closed. The converse implication does not hold as the following example shows.
Example 2.21.
Consider again the Moore family ${\mathcal{F}}$ from Example 2.14 and its closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$. Then one has ()={,$\wp({\mathcal{F}})=\{\,\,\emptyset,\,\{x,z\},\,\{x,w\}\,\,\}$. Note in this context that none of the singleton subsets of $X$ is pseudo-closed because of $\mbox{\rm cl}_{{\mathcal{F}}}(\emptyset)=\{x\}$. The empty set $\emptyset$ is the only inclusion-minimal non-closed set while $\{x,z\}$ and $\{x,w\}$ are additional pseudo-closed sets. Specifically, one has
${\mathbb{G}}_{{\mathcal{F}}}~{}=~{}\{~{}\emptyset\to\{x\},\,\{x,z\}\to\{y\},\,\{x,w\}\to\{y\}~{}\}\,.$
Nevertheless, further simplification is possible. Because of $\emptyset\to\{x\}$, the implication $\{x,w\}\to\{y\}$ is equivalent to $\{w\}\to\{y\}$ (use axioms of functional dependence relations to observe that). To summarize: the simplest implicational generator for the family ${\mathcal{F}}$ seems to be the following list:
$\{~{}\emptyset\to\{x\},\,\{z\}\to\{y\},\,\{w\}\to\{y\}~{}\}\,.$
The well-known result [12, Theorem 7 in § 3.2] says $({\mathbb{G}}_{{\mathcal{F}}})^{\triangleright}={\mathcal{F}}$ for any Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$. In fact, ${\mathbb{G}}_{{\mathcal{F}}}$ is a non-redundant implicational generator of ${\mathcal{F}}$ in sense that the removal of any of its elements leads to a non-generator.
Let us introduce a specific terminology concerning formal implications. Given a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$, a valid implication for ${\mathcal{F}}$ is $(Y\to Z)\in{\mathcal{F}}^{\triangleleft}$. An implication $Y\to Z$ is perfect (for ${\mathcal{F}}$) if $Y\not\in{\mathcal{F}}$ but every strict subset of $Y$ is closed, that is, $Y\in({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$, and $Z=\mbox{\rm cl}_{{\mathcal{F}}}(Y)\setminus Y$. The implication is canonical if it belongs to the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$.
It is immediate that every perfect implication is canonical but the converse does not hold as shown in Example 2.21, where $\{x,z\}\to\{y\}$ is a canonical implication which is not perfect.
Definition 2.22 (implicatively perfect Moore family).
We say that a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$ is implicatively perfect if the set of perfect implications for ${\mathcal{F}}$ is an implicational generator of ${\mathcal{F}}$.
It makes no problem to deduce from non-redundancy of the canonical basis that an equivalent condition is that every pseudo-closed set is an inclusion-minimal non-closed set, that is, $\wp({\mathcal{F}})=({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$. [^6] The concept from Definition 2.22 can be viewed as a counter-part of the concept of a coatomistic lattice in the context of implicational description of the lattice.
Example 2.23.
(semi-graphoids over 3 variables revisited)
Consider again the situation from Example 2.16: $N:=\{a,b,c\}$, $X:={\mathcal{K}}(N)$, and ${\mathcal{F}}:=\mathfrak{sg}\,(N)$. There are six pseudo-closed sets relative to ${\mathcal{F}}$ belonging to one permutational type: the set $\{\,ab|\emptyset,\,ac|b\,\}$ and five of its permutational copies. Each of these sets is minimal in ${\mathcal{P}}(X)\setminus{\mathcal{F}}$; thus, ${\mathcal{F}}$ is implicatively perfect. The canonical basis ${\mathbb{G}}_{{\mathcal{F}}}$ consists of permutational copies of the implication
$\{\,ab|\emptyset,\,ac|b\,\}\to\{\,ac|\emptyset,\,ab|c\,\}\,.$
Thus, the characterization of ${\mathcal{F}}$ in terms of its canonical implicational basis is equivalent to the definition of a semi-graphoid from Section 2.2.2.
The fact that the elements of the canonical basis for semi-graphoids fall down into permutational types is not a surprise because we have the following.
Corollary 3.
If a CI frame $\mathfrak{f}$ is closed under copying and intersection then, for each set $N$, the valid implications $Y\to Z$ and the elements of the canonical implicational basis for ${\mathcal{F}}:=\mathfrak{f}(N)$ fall into permutational equivalence classes.
Proof 2.24.
The arguments are analogous to those in Corollary 2. Any permutation $\pi:N\to N$ defines a simultaneous order automorphism of the posets $(\mathfrak{f}(N),\subseteq)$ and $({\mathcal{P}}({\mathcal{K}}(N)),\subseteq)$. As such an order automorphism of $({\mathcal{F}},\subseteq)$ commutes with the closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$, it preserves the validity of implications.
In the sequel, we freely switch between the description of a CI frame $\mathfrak{f}$ in terms of the families $\mathfrak{f}(N)$ and in terms of the corresponding closure operations.
Definition 2.25 (frame-based closure of a CI model).
Consider a CI frame $\mathfrak{f}$ closed under marginalization, intersection, and satisfying $\mathfrak{f}(N)\subseteq\mathfrak{f}(M)$ if $N\subseteq M$ [^7] As ${\mathcal{K}}(N)\subseteq{\mathcal{K}}(M)$ if $N\subseteq M$, the inclusion $\mathfrak{f}(N)\subseteq\mathfrak{f}(M)$ of set systems makes sense. Given a CI model ${\mathcal{R}}\subseteq{\mathcal{K}}(N)$ over a variable set $N$, the closure of ${\mathcal{R}}$ determined by $\mathfrak{f}(N)$ will shortly be denoted by $\mathfrak{f}({\mathcal{R}}):=\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{R}})$.
The assumptions in Definition 2.25 imply that the definition of $\mathfrak{f}$-closure does not actually depend on $N$: one has $\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{R}})=\mbox{\rm cl}_{\mathfrak{f}(M)}({\mathcal{R}})$ for any $M\supseteq N$. All frames considered in this article satisfy the assumptions of Definition 2.25.
3. Adhesion and self-adhesion
The concept of an adhesion/self-adhesion was introduced by Matúš [23] in the context of polymatroids. As pinpointed by Csirmaz [8], the adhesion concept is a special case of the concept of an amalgam for polymatroids. Note that amalgams were introduced and studied much earlier in matroid theory [25].
It has been observed recently in [2] that these concepts can also be defined within a simpler discrete environment of formal CI structures. We follow this link here. An auxiliary concept of consonant models is needed.
Definition 3.1 (consonant models, amalgam, adhesion).
  • Let $N,M$ be two variable sets, possibly intersecting. Models ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$ will be called consonant if ${\mathcal{M}}^{\downarrow N\cap M}={\mathcal{R}}^{\downarrow N\cap M}$.
  • Given consonant models ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$, their amalgam is any “extended” model ${\mathcal{E}}\subseteq{\mathcal{K}}(NM)$ over $NM$ such that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$ and ${\mathcal{E}}^{\downarrow M}={\mathcal{R}}$.
  • An adhesion of ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$ is their amalgam ${\mathcal{E}}\subseteq{\mathcal{K}}(NM)$ that, moreover, satisfies $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap M)\,\,[{\mathcal{E}}]$ (see Definition 2.9).
Note that any pair of models ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$ is consonant if $|N\cap M|\leq 1$. The consonancy is a necessary condition for the existence of an amalgam. Observe that any pair of consonant models has an adhesion: take the union of ${\mathcal{M}}$, ${\mathcal{R}}$, and the subset of ${\mathcal{K}}(NM)$ delimiting the global CI statement $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap M)$.
Unlike the algebraic operations from Definition 2.2, amalgams and adhesions are not uniquely determined by their input models ${\mathcal{M}}$ and ${\mathcal{R}}$. On the other hand, (set-theoretical) intersection of two amalgams/adhesions of ${\mathcal{M}}$ and ${\mathcal{R}}$ is their amalgam/adhesion; thus, their least amalgam/adhesion exists.
Example 3.2.
Take $N:=\{a,b,c,d\}$, $M:=\{c,d,e\}$, and put :=${\mathcal{M}}:=\{\,ab|\emptyset,\,cd|\emptyset\,\}$, :=${\mathcal{R}}:=\{\,cd|\emptyset,\,cd|e\,\}$, and :=${\mathcal{R}}^{\prime}:=\{\,cd|e\,\}$. Then ${\mathcal{M}}$ and ${\mathcal{R}}$ are consonant models, while ${\mathcal{M}}$ and ${\mathcal{R}}^{\prime}$ are not. The least amalgam of ${\mathcal{M}}$ and ${\mathcal{R}}$ is $\{\,ab|\emptyset,\,cd|\emptyset,cd|e\,\}$ and the least adhesion of ${\mathcal{M}}$ and ${\mathcal{R}}$ is
$\{\,ab|\emptyset,\,cd|\emptyset,\,cd|e,\,\,\,ae|cd,\,ae|bcd,\,be|cd,\,be|acd\,\}\,,$
where the last 4 statements represent the global CI statement $ab\perp\!\!\!\perp e\,|\,cd$.
3.1. Self-adhesion
Our basic notion is self-adhesion of a model ${\mathcal{M}}$ is nothing but the adhesion of ${\mathcal{M}}$ with its own $\varphi$-copy, where $\varphi$ is an appropriate bijection. The derived concept of self-adhesivity is relative to a considered CI frame.
Definition 3.3 (self-adhesion).
  • Let $\mathfrak{f}$ be a CI frame closed under marginalization and copying such that $\mathfrak{f}\subseteq\mathfrak{sg}$. Given ${\mathcal{M}}\in\mathfrak{f}(N)$ and $L\subseteq N$ we say that ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{f}$ if, for every bijection $\varphi:N\to M$ such that $L=N\cap M$ and $\varphi_{\,|L}$ is the identity, an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ exists which is in the family $\mathfrak{f}(NM)$:
    $\forall\,\varphi:N\to M~{}\mbox{bijection with $L=N\cap M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$~{}}~{}~{}\exists\,\,{\mathcal{E}}\in\mathfrak{f}(NM)$
    such that N=M=φ() and ($\mbox{such that~{} ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$,~{} ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}})$ and $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$}\,.$
    Then ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive relative to $\mathfrak{f}$ if it is self-adhesive at any $L\subseteq N$.
  • The symbol $\mathfrak{f}^{\diamond}(N)$ will denote the family of self-adhesive models ${\mathcal{M}}\in\mathfrak{f}(N)$ relative to $\mathfrak{f}$. If $\mathfrak{f}(N)=\mathfrak{f}^{\diamond}(N)$ then we say that $\mathfrak{f}(N)$ is self-adhesive. The frame $\mathfrak{f}$ is self-adhesive if this holds for any variable set $N$.
  • Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $L\subseteq N$, if there exists the least model ${\mathcal{R}}\in\mathfrak{f}(N)$ such that ${\mathcal{M}}\subseteq{\mathcal{R}}$ and ${\mathcal{R}}$ is self-adhesive at $L$ relative to $\mathfrak{f}$ then we will call ${\mathcal{R}}$ the self-adhesive closure of ${\mathcal{M}}$ at $L$ (relative to $\mathfrak{f}$) and denote it by $\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$.
Note that the assumption $\mathfrak{f}\subseteq\mathfrak{sg}$ is to make the adhesion concept sensible. The other two assumptions (on marginalization and copying) are also needed to define the concept of self-adhesion. Clearly, a model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L\subseteq N$ relative to $\mathfrak{f}$ if and only if $\mathfrak{f}^{\diamond}({\mathcal{M}}|L)={\mathcal{M}}$.
Remarks.  We have a few additional related remarks.
  1. (1)
    One can equivalently introduce a self-adhesive model ${\mathcal{M}}$ at a set $L$ by a formally weaker condition:
    $\exists\,\psi:N\to O~{}\mbox{bijection with $L=N\cap O~{}\mbox{and}~{}\psi_{\,|L}=\mbox{\rm id}_{L}$}~{}~{}~{}\exists\,\,\widetilde{{\mathcal{E}}}\in\mathfrak{f}(NO)\,:$
    ~$\widetilde{{\mathcal{E}}}^{\downarrow N}={\mathcal{M}}~{}\&~{}~{}\widetilde{{\mathcal{E}}}^{\downarrow O}=\psi({\mathcal{M}})~{}\&~{}~{}(N\setminus O)\perp\!\!\!\perp(O\setminus N)\,|\,L\,\,[\widetilde{{\mathcal{E}}}]\,.$
    This follows from the assumption that $\mathfrak{f}$ is closed under copying. Having fixed $\psi:N\to O$, for any considered bijection $\varphi:N\to M$ we define $\rho:NO\to NM$ by $\rho(i):=i$ for $i\in N$ and $\rho(i):=\varphi(\psi_{-1}(i))$ for $i\in O\setminus N$. Then ${\mathcal{E}}:=\rho(\widetilde{{\mathcal{E}}})$ is the desired adhesion for $\varphi$. Therefore, the choice of the bijection $\varphi$ is irrelevant and the validity of the self-adhesivity condition is, in fact, an intrinsic property of ${\mathcal{M}}$ (relative to $\mathfrak{f}$).
  2. (2)
    This is to explain why the self-adhesivity concept is relative to a CI frame. Having two frames $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$, imagine one finds a model ${\mathcal{M}}\in\mathfrak{f}_{1}(N)$ and a bijection $\varphi:N\to M$ with $L=N\cap M$ such that a self-adhesion ${\mathcal{E}}\in\mathfrak{f}_{2}(NM)$ of ${\mathcal{M}}$ for $\varphi$ exists while there is no such a self-adhesion within $\mathfrak{f}_{1}(NM)$. Then ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{f}_{2}$ but not relative to $\mathfrak{f}_{1}$. This really can happen as Example 5.4 shows (see Section 5.6).
  3. (3)
    It follows from the definition that, given ${\mathcal{M}}\in\mathfrak{f}(N)$, the question whether ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ is decidable on basis of the characterization/description of families $\mathfrak{f}(M)$, where $|N|\leq|M|\leq 2\cdot|N|$. Nevertheless, further reduction is possible. For example, it is immediate that any ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L:=N$. Additional reductions, based on further assumptions on the frame, are discussed in Section 3.2.
If the considered frame $\mathfrak{f}$ is closed under intersection then the self-adhesive closure of any model at any set is easy to compute by applying the $\mathfrak{f}$-closure. In particular, the question of self-adhesivity testing can be simplified.
Lemma 3.4.
Let $\mathfrak{f}$ be a CI frame closed under marginalization, copying, and intersection such that $\mathfrak{f}\subseteq\mathfrak{sg}$. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $L\subseteq N$, the self-adhesive closure ${\mathcal{R}}=\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$ exists and can be computed as follows. Having fixed a bijection $\varphi:N\to M$ with $L=N\cap M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$, apply the next steps:
  1. (1)
    take the least self-adhesion of ${\mathcal{M}}$: ${\mathcal{E}}:={\mathcal{M}}\cup\varphi({\mathcal{M}})\cup[N\setminus M,M\setminus N|L]$,
  2. (2)
    apply the $\mathfrak{f}$-closure to it: ~$\widetilde{{\mathcal{E}}}:=\mathfrak{f}({\mathcal{E}})$,
  3. (3)
    apply marginalization to it: :=~${\mathcal{R}}:=\widetilde{{\mathcal{E}}}^{\downarrow N}$.
Proof 3.5.
Extend the bijection $\varphi:N\to M$ by the restriction of $\varphi_{-1}$ to $M\setminus N$ to get a permutation $\psi$ of $NM$. Observe that $\psi({\mathcal{E}})={\mathcal{E}}$. Because $\mathfrak{f}$ is closed under copying, by arguments in Corollary 1, the $\mathfrak{f}$-closure over $NM$ commutes with $\psi$ and one even has $-~\psi(\widetilde{{\mathcal{E}}})=\widetilde{{\mathcal{E}}}$. It follows from the construction and the fact that $\mathfrak{f}$ is closed under marginalization that ${\mathcal{M}}\subseteq{\mathcal{R}}\in\mathfrak{f}(N)$. Using that $-~\widetilde{{\mathcal{E}}}$ is fixed by $\psi$ and the fact that copying commutes with marginalization, one has
~$\widetilde{{\mathcal{E}}}^{\downarrow M}=\widetilde{{\mathcal{E}}}^{\downarrow\psi(N)}=\psi(\widetilde{{\mathcal{E}}})^{\downarrow\psi(N)}=\psi_{\,|N}(\widetilde{{\mathcal{E}}}^{\downarrow N})=\varphi(\widetilde{{\mathcal{E}}}^{\downarrow N})=\varphi({\mathcal{R}})\,,$
which implies, by Definition 3.3, that ${\mathcal{R}}$ is self-adhesive at $L$.
Given ${\mathcal{R}}^{\prime}\in\mathfrak{f}(N)$ which is self-adhesive at $L$, and satisfies ${\mathcal{M}}\subseteq{\mathcal{R}}^{\prime}$, it is easy to observe that the respective adhesion ${\mathcal{E}}^{\prime}\in\mathfrak{f}(NM)$ satifies ${\mathcal{E}}\subseteq{\mathcal{E}}^{\prime}$, which yields $\widetilde{{\mathcal{E}}}\subseteq{\mathcal{E}}^{\prime}$ and, hence, ${\mathcal{R}}\subseteq{\mathcal{R}}^{\prime}$. Therefore, one indeed has ${\mathcal{R}}=\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$.
The next example illustrates the method from Lemma 3.4.
Example 3.6.
Take $N:=\{a,b,c\}$ and put :=${\mathcal{M}}:=\{\,ab|c\,\}$; note that ${\mathcal{M}}\in\mathfrak{sg}(N)$. We are going to test self-adhesivity of ${\mathcal{M}}$ at the set $L:=\{b,c\}$ relative to the semi-graphoidal frame $\mathfrak{sg}$. As explained (in the remark n. 1 above) it suffices to fix a variable set $M:=\{\bar{a},b,c\}$ and a bijection $\varphi:N\to M$ defined by $\varphi(a):=\bar{a}$, $\varphi(b):=b$, and $\varphi(c):=c$, (see Figure 2) and to find an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ which belongs to $\mathfrak{sg}(NM)$. To this end we consider their least adhesion =${\mathcal{E}}=\{\,ab|c,\,\bar{a}b|c,\ a\bar{a}|bc\,\}$ and take its semi-graphoidal closure
$\mathfrak{sg}({\mathcal{E}})\,=\,\{\,ab|c,\,ab|\bar{a}c,\,a\bar{a}|c,\,a\bar{a}|bc,\,\bar{a}b|c,\,\bar{a}b|ac\,\}\,.$
Since $\mathfrak{sg}({\mathcal{E}})^{\downarrow N}={\mathcal{M}}$, one has $\mathfrak{sg}^{\diamond}({\mathcal{M}}|L)={\mathcal{M}}$, and ${\mathcal{M}}$ is self-adhesive at $L$ relative to $\mathfrak{sg}$. Observe that one has $\mathfrak{sg}({\mathcal{E}})^{\downarrow M}=\{\,\bar{a}b|c\,\}=\varphi({\mathcal{M}})$.
Figure 2. The bijection $\varphi:N\to M$ considered in Example 3.6.
This was an example of self-adhesivity at a set. On the other hand, Example 5.1 (in Section 5.3) shows that it may happen that $\mathfrak{sg}^{\diamond}({\mathcal{M}}|L)\neq{\mathcal{M}}$.
Lemma 3.7.
The discrete probabilistic CI frame $\mathfrak{pr}$ is self-adhesive: $\mathfrak{pr}^{\diamond}=\mathfrak{pr}$.
Proof 3.8.
[…]
3.2. Lifting and tight replication
This section is to show that, if a frame is closed under appropriate operations then the involved models are self-adhesive at sets of particular cardinality.
Definition 3.9 (lifting, tight replication).
  • Let $N\subseteq O$ be two variable sets in inclusion and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. The lift of ${\mathcal{M}}$ to $O$ is the following model over $O$:
    ${\mathcal{M}}_{N\nearrow O}~{}:=~{}\{\,ij|K\in{\mathcal{K}}(O)\,:\ ij\cap(O\setminus N)\neq\emptyset~{}~{}\mbox{or}~{}~{}ij|(K\cap N)\in{\mathcal{M}}\,\}\,.$
    A CI frame $\mathfrak{f}$ is closed under lifting, if for any ${\mathcal{M}}\in\mathfrak{f}(N)$ and any superset $O\supseteq N$, one has ${\mathcal{M}}_{N\nearrow O}\in\mathfrak{f}(O)$.
  • Let $N,M$ be two variable sets of cardinality $n\geq 1$ such that $|N\cap\/M|=n-1$. Let us denote $L:=N\cap M$, $u:=N\setminus L$, $v:=M\setminus L$, and introduce a mapping $\rho:NM\to N$ by $\rho(i):=i$ for $i\in N$ and $\rho(v):=u$. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$, its tight replication (based on $\rho$) is the model ${\mathcal{M}}_{v\|u}\subseteq{\mathcal{K}}(NM)$ defined as follows:
    $\mbox{given $ij|K\in{\mathcal{K}}(NM)$,}\quad ij|K\in{\mathcal{M}}_{v\|u}~{}~{}\mbox{if and only if}~{}~{}\rho(i)\perp\!\!\!\perp\rho(j)\,|\,\rho(K)\,\,[{\mathcal{M}}],$
    under conventions that $a\perp\!\!\!\perp b\,|\,C\,\,[{\mathcal{M}}]$ whenever $ab\cap C\neq\emptyset$
    and $u\perp\!\!\!\perp u\,|\,C\,\,[{\mathcal{M}}]$ for $C\subseteq L$ means $u\perp\!\!\!\perp(L\setminus C)\,|\,C\,\,[{\mathcal{M}}]$
    A CI frame $\mathfrak{f}$ is closed under tight replication, if for any such a pair $N,M$ of variable sets and ${\mathcal{M}}\in\mathfrak{f}(N)$, one has ${\mathcal{M}}_{v\|u}\in\mathfrak{f}(NM)$.
Lifting complies with marginalization in the sense that, for every ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $N\subseteq O$, one has $({\mathcal{M}}_{N\nearrow O})^{\downarrow N}={\mathcal{M}}$. Analogously, the tight replication also complies with marginalization in the sense that $({\mathcal{M}}_{v\|u})^{\downarrow N}={\mathcal{M}}$.
Figure 3. The mapping $\rho:NM\to N$ considered in Example 3.10.
Example 3.10.
Put $N:=\{a,b\}$, $O:=\{a,b,c\}$, ${\mathcal{M}}:=\emptyset$, and :=${\mathcal{R}}:=\{\,ab|\emptyset\,\}$. Then ${\mathcal{M}}_{N\nearrow O}=\{\,ac|\emptyset,\,ac|b,\,bc|\emptyset,\,bc|a\,\}$. The interpretation of the lifting is that each element of $O\setminus N$ (in this case, only the variable $c$) becomes independent of the rest. Thus, in case of ${\mathcal{R}}$, one gets the full model: ${\mathcal{R}}_{N\nearrow O}={\mathcal{K}}(O)$.
The interpretation of the tight replication is that one extends the model by introducing a new variable $v$ which is an identical copy of one of the original variables, namely of $u$. To illustrate this concept consider the situation when one has $N:=\{a,b,c\}$, $M:=\{b,c,d\}$, $L:=\{b,c\}$, $u:=a$ and $v:=d$. Define $\rho(a):=a$, $\rho(b):=b$, $\rho(c):=c$, $\rho(d):=a$ (see Figure 3 for illustration) and put ${\mathcal{M}}:=\emptyset$ and :=${\mathcal{R}}:=\{\,ab|c\,\}$. Then
${\mathcal{M}}_{v\|u}=\{\,\,ab|d,\,ab|cd,\,ac|d,\,ac|bd,\,bd|a,\,bd|ac,\,cd|a,\,cd|ab,\,ad|bc\,\,\}\,,$
where the first 8 statements are there because of the first convention saying that $a\perp\!\!\!\perp b\,|\,C\,\,[{\mathcal{M}}]$ whenever $ab\cap C\neq\emptyset$ and the last statement $ad|bc$ is added because of the second convention $u\perp\!\!\!\perp u\,|\,C\,\,[{\mathcal{M}}]$ if and only if $u\perp\!\!\!\perp(L\setminus C)\,|\,C\,\,[{\mathcal{M}}]$. In case of ${\mathcal{R}}$ one gets ${\mathcal{R}}_{v\|u}={\mathcal{M}}_{v\|u}\cup\{\,ab|c,\,bd|c,\,ad|c\,\}$, where $bd|c$ is mapped by $\rho$ to $ab|c$ and $ad|c$ is added because of the second convention.
Remark on terminology.   Operations with supermodular set functions, which are analogous to those in Definition 3.9, were discussed in [30, § 7.1 and § 10.2]. Moreover, the concept of tight replication of CI models is a counterpart of an analogous notion of parallel extension of matroids [25, § 7.2] and of polymatroids [23, § 2]. Nevertheless, the parallel extension of polymatroids has two slightly different reflections in the context of induced CI models. In the case that the considered polymatroid $h$ is tight [8, § 1.3] at the variable $u\in N$, which means that $h(N)=h(N\setminus u)$, it leads to our second concept in Definition 3.9. If, however, $h$ is not tight at $u$ then it leads to the concept of parallel extension of CI models introduced in 2004 by Matúš [22, end of § 4]. The extension we need for our purposes is the one that corresponds to a tight polymatroid. Therefore, to avoid a terminological clash, we have chosen a different term than parallel extension. Our terminology also reflects the crucial role of tightness.
Lemma 3.11.
The CI frames $\mathfrak{pr}$, $\mathfrak{st}$, and $\mathfrak{sg}$ are closed under lifting.
Proof 3.12.
[…]
Given a supermodular function $r:{\mathcal{P}}(N)\to{\mathbb{R}}$ and $N\subseteq O$, one can extend it to a supermodular function $m$ on ${\mathcal{P}}(O)$ by the formula $m(S):=r(S\cap N)$ for any $S\subseteq O$. It is easy to observe that ${\mathcal{M}}^{m}=({\mathcal{M}}^{r})_{N\nearrow O}$.
Given a semi-graphoid ${\mathcal{M}}\in\mathfrak{sg}(N)$ over $N$ and $N\subseteq O$, it is straightforward to observe that ${\mathcal{M}}_{N\nearrow O}$ is a semi-graphoid over $O$.
Lemma 3.13.
The CI frames $\mathfrak{st}$ and $\mathfrak{sg}$ are closed under tight replication.
Proof 3.14.
Consider variable sets $N,M$ assumed in the definition of tight replication and keep the corresponding notational conventions from Definition 3.9. Assume that ${\mathcal{M}}\in\mathfrak{st}(N)$ is induced by a supermodular function $r:{\mathcal{P}}(N)\to{\mathbb{R}}$. Slight modification $\hat{r}$ of $r$ defined by
$\hat{r}(S)~{}:=~{}\left\{\begin{array}[]{cl}r(S)-r(N)+r(L)&\mbox{if $u\in S$,}\\ r(S)&\mbox{if $u\not\in S$,}\\ \end{array}\right.\qquad\mbox{for $S\subseteq N$,}$
is another supermodular function on ${\mathcal{P}}(N)$ which also induces ${\mathcal{M}}$ but, moreover, satisfies $\hat{r}(N)=\hat{r}(L)$. The next step is to define
$m(S)~{}:=~{}\hat{r}(\,(S\cap N)\cup\rho(S\cap M)\,)\qquad\mbox{for~{} $S\subseteq NM$}\,.$
For any $ij|K\in{\mathcal{K}}(NM)$ one has $\Delta m(ij|K)=\Delta\hat{r}\,\langle\rho(i),\rho(j)|\rho(K)\rangle$. The latter expression is non-negative in case $\rho(ij)|\rho(K)\in{\mathcal{K}}(N)$ by supermodularity of $\hat{r}$. Moreover, the expression vanishes if $\rho(ij)\cap\rho(K)\neq\emptyset$. In the remaining case $\rho(i)=u=\rho(j)$ and $C:=\rho(K)\subseteq L$ one has
$\Delta\hat{r}\,\langle\rho(i),\rho(j)|\rho(K)\rangle=\Delta\hat{r}\,\langle u,u|C\rangle=\hat{r}(C)-\hat{r}(uC)=\Delta\hat{r}\,\langle u,L\setminus C|C\rangle\geq 0\,,$
by supermodularity of $\hat{r}$. Hence, $m$ is supermodular and it follows from the definition of tight replication that ${\mathcal{M}}^{m}={\mathcal{M}}_{v\|u}$.
Given ${\mathcal{M}}\in\mathfrak{sg}(N)$, it is enough to show that ${\mathcal{E}}:={\mathcal{M}}_{v\|u}$ is a semi-graphoid. Consider pairwise disjoint sets $i,j,\ell,K\subseteq NM$ and the goal is to verify
$ij|K,\,i\ell|jK\in{\mathcal{E}}~{}\Leftrightarrow~{}i\ell|K,\,ij|\ell K\in{\mathcal{E}}\,.$
One can distinguish several cases determined by the mutual position of variables $u$ and $v$ relative to $i,j,\ell,K$. If $uv\setminus ij\ell K\neq\emptyset$ then the conclusion follows from the facts that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$ and ${\mathcal{E}}^{\downarrow M}$ is a copy of ${\mathcal{M}}$. The same principle applies if $u,v\in K$. In cases $|uv\cap K|=1=|uv\cap ij\ell|$ the first convention in the definition of ${\mathcal{M}}_{v\|u}$ is applied. This also extends to the case $uv=j\ell$. In the cases $uv=ij$ and $uv=i\ell$ one additionally applies the second convention in the definition of ${\mathcal{M}}_{v\|u}$ together with the assumption that ${\mathcal{M}}$ is a semi-graphoid.
The lifting operation is related to adhesions at small sets.
Lemma 3.15.
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame which is closed under marginalization, intersection, and lifting. Then, for every ${\mathcal{M}}\in\mathfrak{f}(N)$ and ${\mathcal{R}}\in\mathfrak{f}(M)$ such that $|N\cap\/M|\leq 1$, an adhesion ${\mathcal{E}}\in\mathfrak{f}(NM)$ of ${\mathcal{M}}$ and ${\mathcal{R}}$ exists.
Proof 3.16.
Put ^$\hat{{\mathcal{M}}}:={\mathcal{M}}_{N\nearrow NM}$ and observe that $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap\/M)\,\,[\hat{{\mathcal{M}}}]$ and ^$\hat{{\mathcal{M}}}^{\downarrow N}={\mathcal{M}}$. Moreover, $|N\cap\/M|\leq 1$ implies ^$\hat{{\mathcal{M}}}^{\downarrow M}={\mathcal{K}}(M)$ and the fact ^$\hat{{\mathcal{M}}}\in\mathfrak{f}(NM)$ follows from the assumption that $\mathfrak{f}$ is closed under lifting. Putting ^$\hat{{\mathcal{R}}}:={\mathcal{R}}_{M\nearrow NM}\in\mathfrak{f}(NM)$ gives $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,(N\cap\/M)\,\,[\hat{{\mathcal{R}}}]$ with ^$\hat{{\mathcal{R}}}^{\downarrow N}={\mathcal{K}}(N)$ and ^$\hat{{\mathcal{R}}}^{\downarrow M}={\mathcal{R}}$. Since $\mathfrak{f}$ is closed under intersection, $-^{\mathcal{E}}:=\hat{{\mathcal{M}}}\cap\hat{{\mathcal{R}}}$ belongs to $\mathfrak{f}(NM)$ and is an adhesion of ${\mathcal{M}}$ and ${\mathcal{R}}$.
Of course, the above observation on adhesivity applies to self-adhesivity.
Corollary 4.
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame which is closed under marginalization, copying, intersection, and lifting. Then, for any $N$, any model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at any set of cardinality 00 or $1$.
Proof 3.17.
Apply Lemma 3.15 to ${\mathcal{M}}$ and ${\mathcal{R}}:=\varphi({\mathcal{M}})$ for a bijection $\varphi:N\to M$.
The tight replication is related to self-adhesions at large sets.
Lemma 3.18.
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame which is closed under marginalization, copying, and tight replication. Then, for any variable set $N$ of cardinality $n\geq 1$, any model ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at any set of cardinality $n-1$.
Proof 3.19.
Consider ${\mathcal{M}}\in\mathfrak{f}(N)$, a subset $L\subseteq N$ with $|L|=n-1$, and a bijection $\varphi:N\to M$ such that $N\cap\/M=L$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$. Put $u:=N\setminus L$ and $v:=M\setminus L$. By the assumption, ${\mathcal{E}}:={\mathcal{M}}_{v\|u}$ belongs to $\mathfrak{f}(NM)$. The definition of tight replication allows one to observe that ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}})$, and $u\perp\!\!\!\perp v\,|\,L\,\,[{\mathcal{E}}]$. Hence, ${\mathcal{E}}$ is an adhesion of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$.
Corollary 5.
If a CI frame $\mathfrak{f}\subseteq\mathfrak{sg}$ is closed under marginalization, copying, intersection, lifting, and tight replication and $|N|\leq 3$ then $\mathfrak{f}^{\diamond}(N)=\mathfrak{f}(N)$. In particular, $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}(N)$ and $\mathfrak{st}^{\diamond}(N)=\mathfrak{st}(N)$ in case $|N|\leq 3$.
Proof 3.20.
Combine Corollary 4, Lemma 3.18 and the fact that any ${\mathcal{M}}\in\mathfrak{f}(N)$ is self-adhesive at $L:=N$. Use Lemmas 2.6, 2.11, 3.11, and 3.13 for the second part.
The second part of Corollary 5 is not surprising in light of Lemma 3.7 since $\mathfrak{sg}(N)=\mathfrak{st}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$. On the other hand, the observations in Section 3.2 allow one to simplify testing self-adhesive models in case $|N|\geq 4$.
4. Iterations of the self-adhesion operator
Observe that copying commutes with marginalization: given ${\mathcal{M}}\in{\mathcal{K}}(N)$, a set $L\subseteq N$, and a bijection $\varphi:N\to M$, one has $\varphi({\mathcal{M}})^{\downarrow\varphi(L)}=\varphi_{|L}({\mathcal{M}}^{\downarrow L})$. Analogously, intersection commutes both with copying and marginalization: given ${\mathcal{M}},{\mathcal{R}}\in{\mathcal{K}}(N)$, a bijection $\varphi:N\to M$ and a set $L\subseteq N$, one has $\varphi({\mathcal{M}}\cap{\mathcal{R}})=\varphi({\mathcal{M}})\cap\varphi({\mathcal{R}})$ and $({\mathcal{M}}\cap{\mathcal{R}})^{\downarrow L}={\mathcal{M}}^{\downarrow L}\cap{\mathcal{R}}^{\downarrow L}$.
Lemma 4.1.
Let $\mathfrak{f}\subseteq\mathfrak{sg}$ be a CI frame closed under copying and marginalization. Then $N\mapsto\mathfrak{f}^{\diamond}(N)$ defines a frame closed under copying and marginalization. If $\mathfrak{f}$ is closed under intersection then $\mathfrak{f}^{\diamond}$ is closed under intersection as well. If $\mathfrak{f}$ is closed under lifting then $\mathfrak{f}^{\diamond}$ is closed under lifting, too.
Proof 4.2.
It is easy to see that ${\mathcal{K}}(N)\in\mathfrak{f}^{\diamond}(N)$ because ${\mathcal{K}}(O)\in\mathfrak{f}(O)$ for any $O\supseteq N$. Hence, $N\mapsto\mathfrak{f}^{\diamond}(N)$ is a frame in sense of Definition 2.1.
Consider ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and a bijection $\varphi:N\to M$. The goal is to show that ${\mathcal{R}}:=\varphi({\mathcal{M}})\in\mathfrak{f}^{\diamond}(M)$. Take $L\subseteq M$ and choose a bijection $\psi:M\to O$ such that $L=M\cap O$, $\psi_{|L}=\mbox{\rm id}_{L}$, and $(O\setminus L)\cap N=\emptyset$. By the remark n. 1 below Definition 3.3, it is enough to show that ${\mathcal{R}}$ and $\psi({\mathcal{R}})$ have an adhesion in $\mathfrak{f}(MO)$. Put $K:=\varphi_{-1}(L)\cup(O\setminus L)$ and define a bijection $\rho:N\to K$ by $\rho(i):=i$ for $i\in\varphi_{-1}(L)$ and $\rho(i):=\psi(\varphi(i))$ for $i\in N\setminus\varphi_{-1}(L)$. Since ${\mathcal{M}}$ is self-adhesive at $\varphi_{-1}(L)$, an extended model ${\mathcal{E}}\in\mathfrak{f}(NK)$ exists with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow K}=\rho({\mathcal{M}})$ and $(N\setminus\varphi_{-1}(L))\perp\!\!\!\perp(K\setminus\varphi_{-1}(L))\,|\,\varphi_{-1}(L)\,\,[{\mathcal{E}}]$. Extend $\varphi$ by putting $\varphi(i):=i$ for $i\in K\setminus N=O\setminus L$ and have $\varphi:NK\to MO$. Then $\varphi({\mathcal{E}})\in\mathfrak{f}(MO)$ by the assumption on $\mathfrak{f}$. The fact that copying commutes with marginalization together with the identity $\varphi_{|K}\circ\rho=\psi\circ\varphi_{|N}$ allows one to observe $\varphi({\mathcal{E}})^{\downarrow M}={\mathcal{R}}$, $\varphi({\mathcal{E}})^{\downarrow O}=\psi({\mathcal{R}})$ and $(M\setminus L)\perp\!\!\!\perp(O\setminus L)\,|\,L\,\,[\,\varphi({\mathcal{E}})\,]$.
Consider ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and $M\subseteq N$. To show ${\mathcal{M}}^{\downarrow M}\in\mathfrak{f}^{\diamond}(M)$ take $L\subseteq M$ and choose a bijection $\psi:M\to O$ with $L=N\cap O$ and $\psi_{|L}=\mbox{\rm id}_{L}$. Again, it is enough to show that ${\mathcal{M}}^{\downarrow M}$ and $\psi({\mathcal{M}}^{\downarrow M})$ have an adhesion in $\mathfrak{f}(MO)$. Choose a set $K$ with $K\cap NO=\emptyset$ and extend $\psi$ to a bijection from $N$ to $OK$. Since ${\mathcal{M}}$ is self-adhesive at $L$, a model ${\mathcal{E}}\in\mathfrak{f}(NOK)$ exists with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow OK}=\psi({\mathcal{M}})$ and $(N\setminus L)\perp\!\!\!\perp(OK\setminus L)\,|\,L\,\,[{\mathcal{E}}]$. The assumption on $\mathfrak{f}$ gives ${\mathcal{E}}^{\downarrow MO}\in\mathfrak{f}(MO)$. Clearly, $({\mathcal{E}}^{\downarrow MO})^{\downarrow M}={\mathcal{E}}^{\downarrow M}=({\mathcal{E}}^{\downarrow N})^{\downarrow M}={\mathcal{M}}^{\downarrow M}$, and the fact that copying commutes with marginalization allows one to observe $({\mathcal{E}}^{\downarrow MO})^{\downarrow O}=\psi({\mathcal{M}}^{\downarrow M})$. It is immediate that $(M\setminus L)\perp\!\!\!\perp(O\setminus L)\,|\,L\,\,[\,{\mathcal{E}}^{\downarrow MO}\,]$.
Assume ${\mathcal{M}},{\mathcal{R}}\in\mathfrak{f}^{\diamond}(N)$ and $\varphi:N\to M$ with $L:=N\cap\,M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$. To show that ${\mathcal{M}}\cap{\mathcal{R}}$ is self-adhesive at $L$ one finds the respective extensions ${\mathcal{M}}\mapsto{\mathcal{E}}_{{\mathcal{M}}}$, ${\mathcal{R}}\mapsto{\mathcal{E}}_{{\mathcal{R}}}$, both in $\mathfrak{f}(NM)$ and puts ${\mathcal{E}}:={\mathcal{E}}_{{\mathcal{M}}}\cap{\mathcal{E}}_{{\mathcal{R}}}\in\mathfrak{f}(NM)$, by the assumption on $\mathfrak{f}$. Since the intersection commutes both with marginalization and copying one gets ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}\cap{\mathcal{R}}$ and ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}}\cap{\mathcal{R}})$. The definition of ${\mathcal{E}}$ also implies $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$.
Assume ${\mathcal{M}}\in\mathfrak{f}^{\diamond}(N)$ and, without loss of generality, $O\supseteq N$ with $O\setminus N$ being a singleton $o$. We distinguish two cases. The first one is to show that ${\mathcal{M}}_{N\nearrow O}$ is self-adhesive at $L\subseteq N$ relative to $\mathfrak{f}$. Consider a bijection $\varphi:O\to\tilde{M}$ with $L=O\cap\tilde{M}$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$; put $M:=\tilde{M}\setminus\varphi(o)$. The assumption on ${\mathcal{M}}$ implies the existence of ${\mathcal{E}}\in\mathfrak{f}(NM)$ with ${\mathcal{E}}^{\downarrow N}={\mathcal{M}}$, ${\mathcal{E}}^{\downarrow M}=\varphi({\mathcal{M}})$ and $(N\setminus M)\perp\!\!\!\perp(M\setminus N)\,|\,L\,\,[{\mathcal{E}}]$. Then $O\tilde{M}$ is a disjoint union of $NM$ and $o\varphi(o)$ and we can put ${\mathcal{R}}:={\mathcal{E}}_{NM\nearrow O\tilde{M}}$, which belongs to $\mathfrak{f}(O\tilde{M})$, by the assumption on $\mathfrak{f}$. The definition of lifting and the fact that it commutes with copying allow one to observe ${\mathcal{R}}^{\downarrow O}={\mathcal{M}}_{N\nearrow O}$, ${\mathcal{R}}^{\downarrow\tilde{M}}=\varphi({\mathcal{M}}_{N\nearrow O})$ and $(O\setminus\tilde{M})\perp\!\!\!\perp(\tilde{M}\setminus O)\,|\,L\,\,[{\mathcal{R}}]$.
The second case is to show that ${\mathcal{M}}_{N\nearrow O}$ is self-adhesive at $Lo$ (for $L\subseteq N$) relative to $\mathfrak{f}$. Consider a bijection $\varphi:O\to\tilde{M}$ with $Lo:=O\cap\tilde{M}$ and $\varphi$ identical on $Lo$ ; put $M:=\tilde{M}\setminus o$. Choose an adhesion ${\mathcal{E}}\in\mathfrak{f}(NM)$ of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ (at $L$). As $O\tilde{M}=NMo$ put ${\mathcal{R}}:={\mathcal{E}}_{NM\nearrow O\tilde{M}}\in\mathfrak{f}(O\tilde{M})$. Analogous arguments allow one to show that ${\mathcal{R}}$ is an adhesion of ${\mathcal{M}}_{N\nearrow O}$ and $\varphi({\mathcal{M}}_{N\nearrow O})$.
Note that the duality mapping commutes with copying: given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and a bijection $\varphi:N\to M$, one has $\varphi({\mathcal{M}})^{\star}=\varphi({\mathcal{M}}^{\star})$. On the other hand, it does not commute with marginalization: take $N:=\{a,b,c\}$, $M:=\{a,b\}$, and :=${\mathcal{M}}:=\{\,ab|\emptyset\,\}$, in which case $({\mathcal{M}}^{\downarrow M})^{\star}={\mathcal{M}}$ while $({\mathcal{M}}^{\star})^{\downarrow M}=\emptyset$. Therefore, a natural question arised whether, given a CI frame $\mathfrak{f}$ closed under duality, the frame $\mathfrak{f}^{\diamond}$ is closed under duality too. The answer is no: both $\mathfrak{sg}^{\diamond}$ (Example 5.2) and $\mathfrak{st}^{\diamond}$ (Example 5.3) are not closed under duality. An analogous question concerning the operation of tight replication remains open.
{open}
Find out whether $\mathfrak{sg}$ or $\mathfrak{st}$ is an example of a CI frame $\mathfrak{f}\subseteq\mathfrak{sg}$ closed under tight replication such that $\mathfrak{f}^{\diamond}$ is not closed under tight replication.
Thus, we know from Lemma 4.1 that, if $\mathfrak{f}$ is a CI frame allowing to introduce self-adhesive models, then $\mathfrak{f}^{\diamond}$ is another such a frame.
Lemma 4.3.
The self-adhesivity operator $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$, applied to CI frames $\mathfrak{f}\subseteq\mathfrak{sg}$ closed under copying and marginalization, is
  • anti-extensive: $\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$, and
  • isotone: $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}$ implies $\mathfrak{f}^{\diamond}_{1}\subseteq\mathfrak{f}^{\diamond}_{2}$.
Proof 4.4.
This is easy to observe on basis of Definition 3.3.
Corollary 6.
Let $\mathfrak{f}$ be a CI frame closed under copying and marginalization such that $\mathfrak{pr}\subseteq\mathfrak{f}\subseteq\mathfrak{sg}$. Then $\mathfrak{f}^{\diamond}$ has the same properties and $\mathfrak{pr}\subseteq\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$.
Proof 4.5.
Use the first claim in Lemma 4.1. Then Lemmas 2.3, 3.7 and 4.3 allow one to write $\mathfrak{pr}=\mathfrak{pr}^{\diamond}\subseteq\mathfrak{f}^{\diamond}\subseteq\mathfrak{f}$.
Thus, the new CI frame $\mathfrak{f}^{\diamond}$ yields a tighter approximation of $\mathfrak{pr}$ than the original frame $\mathfrak{f}$. If, moreover, $\mathfrak{f}$ is closed under intersection, then, by the second claim in Lemma 4.1, $\mathfrak{f}^{\diamond}$ is also closed under intersection. Therefore, by facts from Section 2.3, for any $N$, the family $\mathfrak{f}^{\diamond}(N)$ can be characterized in terms of CI implications (in the form of Horn clauses), which break down into permutational equivalence classes. By Corollary 6, each of these $\mathfrak{f}^{\diamond}$-rules
  • is a valid probabilistic CI implication, while
  • any of the $\mathfrak{f}$-rules (and, thus, any semi-graphoidal rule) is a valid $\mathfrak{f}^{\diamond}$-rule.
Additionally, if $\mathfrak{f}$ is closed under intersection then, by utilizing the algorithm in Lemma 3.4 as a subroutine, we can compute the $\mathfrak{f}^{\diamond}$-closure of any model (see Definition 2.25).
Lemma 4.6.
Let $\mathfrak{f}$ be a CI frame closed under marginalization, copying, and intersection such that $\mathfrak{f}\subseteq\mathfrak{sg}$ and $\mathfrak{f}^{\diamond}(N)\subseteq\mathfrak{f}^{\diamond}(M)$ if $N\subseteq M$. Then the $\mathfrak{f}^{\diamond}$-closure of any given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ can be computed by the following procedure:
  1. (1)
    start with ${\mathcal{M}}_{0}:={\mathcal{M}}$,
  2. (2)
    iteratively compute ${\mathcal{M}}_{i+1}:=\bigcup_{L\subseteq N}\mathfrak{f}^{\diamond}({\mathcal{M}}_{i}|L)$ for $i=0,1,\dots$,
  3. (3)
    if ${\mathcal{M}}_{i+1}={\mathcal{M}}_{i}$, then output $\mathfrak{f}^{\diamond}({\mathcal{M}}):={\mathcal{M}}_{i}$ and stop.
Proof 4.7.
An ascending chain of CI models ${\mathcal{M}}_{0}\subseteq\cdots\subseteq{\mathcal{M}}_{i}\subseteq\cdots\subseteq{\mathcal{K}}(N)$ is obtained by the iteration, which must eventually stabilize, say at ${\mathcal{M}}_{k}$. Since $\mathfrak{f}^{\diamond}({\mathcal{M}})\in\mathfrak{f}(N)$ is self-adhesive at any $L\subseteq N$ relative to $\mathfrak{f}$, ${\mathcal{M}}_{i}\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$ implies, by Definition 3.3, $\mathfrak{f}^{\diamond}({\mathcal{M}}_{i}|L)\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$. Thus, one can show by induction that ${\mathcal{M}}_{i}\subseteq\mathfrak{f}^{\diamond}({\mathcal{M}})$ at any stage of the algorithm. Since ${\mathcal{M}}_{k}$ is a fixed point of the map ${\mathcal{M}}\mapsto\bigcup_{L\subseteq N}\mathfrak{f}^{\diamond}({\mathcal{M}}|L)$, it is self-adhesive relative to $\mathfrak{f}$ at any subset $L\subseteq N$ (simultaneously), and, thus, it equals to the $\mathfrak{f}^{\diamond}$-closure of ${\mathcal{M}}$.
We wish to know whether or not the self-adhesivity operator $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$ is idempotent, which means $\mathfrak{f}^{\diamond\diamond}=\mathfrak{f}^{\diamond}$. [^8] By Lemma 4.3, $\mathfrak{f}\mapsto\mathfrak{f}^{\diamond}$ would then be interpretable as the “interior” operation, which is a kind of counterpart of the “closure” operation discussed in Section 2.3. Since $\mathfrak{sg}$ is closed under intersection, we were able to use a refined version of the procedure from Lemma 4.6 to compute the iterated self-adhesive models $\mathfrak{sg}^{\diamond\diamond}(N)$ in case $|N|=4$. The basic idea of the algorithm is described in Appendix B. The result of our long computations was the conclusion that $\mathfrak{sg}^{\diamond\diamond}(N)=\mathfrak{sg}^{\diamond}(N)$ if $|N|=4$. An analogous question concerning the $\mathfrak{st}$-frame remains open.
{open}
Find out whether ${\mathcal{M}}\in\mathfrak{st}^{\diamond}(N)\setminus\mathfrak{st}^{\diamond\diamond}(N)$ exists in case $|N|=4$.
5. The case of a small number of variables
In this section we present the results of our computations. We also give a few instructive examples obtained as their results. A summary of numerical data about the lattices we consider in case $|N|=4$ is available in Appendix C.
5.1. Semi-graphoids over 4 variables
Let us start our overview by discussing the family of semi-graphoids $\mathfrak{sg}(N)$ in case $|N|=4$. One has 26 4242642426\,42426 424 semi-graphoids over any 4-element variable set and they form 1512 permutational equivalence classes. The lattice $(\mathfrak{sg}(N),\subseteq)$ has in this case 181 (meet) irreducible elements which fall into 20 permutational types. On the other hand, it only has 37 coatoms falling down into 10 types; thus, the lattice is not coatomistic. In particular, the description of $(\mathfrak{sg}(N),\subseteq)$ in terms (of types) of irreducible models is relatively complicated.
On the other hand, the description of the lattice $(\mathfrak{sg}(N),\subseteq)$ in terms of implicational generators is very easy. The canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$ for ${\mathcal{F}}=\mathfrak{sg}(N)$ in case $|N|=4$ consists of two permutational types only, namely
$\{\,ij|\emptyset,\,i\ell|j\,\}\to\{\,i\ell|\emptyset,\,ij|\ell\,\}~{}~{}\mbox{and}~{}~{}\{\,ij|k,\,i\ell|jk\,\}\to\{\,i\ell|k,\,ij|k\ell\,\}\,.$
Thus, the lattice $(\mathfrak{sg}(N),\subseteq)$ is implicatively perfect. This is not a surprise because the semi-graphoidal frame has, by its definition, a finite “axiomatic” characterization. Informally, by such a characterization of a frame $\mathfrak{f}$ we mean that there is a finite number of schemes which, for each $|N|$, yield an implicational generator for the $\mathfrak{f}$-closure operation.
5.2. Structural semi-graphoids over 4 variables
The family $\mathfrak{st}(N)$ in case $|N|=4$ contains 22 1082210822\,10822 108 models, which fall down into $1285$ permutational equivalence classes. The lattice $(\mathfrak{st}(N),\subseteq)$ is known to be coatomistic for any $N$. Moreover, in the discussed case $|N|=4$, it has the same 37 coatoms as the semi-graphoidal lattice $(\mathfrak{sg}(N),\subseteq)$. Thus, the description of the family $\mathfrak{st}(N)$ in terms of irreducible models is much easier in this case than the description of $\mathfrak{sg}(N)$ in these terms.
As concerns the description of the lattice $(\mathfrak{st}(N),\subseteq)$ in terms of implicational generators, we computed the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$ for ${\mathcal{F}}=\mathfrak{st}(N)$ in case $|N|=4$. What we found out is that it decomposes into 7 permutational types. More specifically, it contains two semi-graphoidal types mentioned in Section 5.1 and five additional permutational types listed below:
  • (E:1)
    $\{\,ij|k,\,ik|\ell,\,i\ell|j\,\}\,\to\,\{\,ij|\ell,\,ik|j,\,i\ell|k\,\}$,
  • (E:2)
    $\{\,ij|k,\,i\ell|j,\,jk|\ell,\,k\ell|i\,\}\,\to\,\{\,ij|\ell,\,i\ell|k,\,jk|i,\,k\ell|j\,\}$,
  • (E:3)
    $\{\,ij|k\ell,\,ik|\emptyset,\,j\ell|\emptyset,\,k\ell|ij\,\}\,\to\,\{\,ij|\emptyset,\,ik|j\ell,\,j\ell|ik,\,k\ell|\emptyset\,\}$,
  • (E:4)
    $\{\,ij|\emptyset,\,ij|k\ell,\,k\ell|i,\,k\ell|j\,\}\,\to\,\{\,ij|k,\,ij|\ell,\,k\ell|\emptyset,\,k\ell|ij\,\}$,
  • (E:5)
    $\{\,ij|k,\,jk|i\ell,\,i\ell|j,\,k\ell|\emptyset\,\}\,\to\,\{\,ij|k\ell,\,jk|i,\,i\ell|\emptyset,\,k\ell|j\,\}$.
Note that our results agree with the analysis presented in [31, § V.A], or with the rules A3–A7 from [28]. In particular, the lattice $(\mathfrak{st}(N),\subseteq)$ is implicatively perfect in case $|N|=4$. Moreover, both the semi-graphoidal rules and the rules (E:1)–(E:5) can be generalized to any higher cardinality than $|N|\geq 4$ [6, 5].
5.3. Self-adhesive semi-graphoids over 4 variables
The family $\mathfrak{sg}^{\diamond}(N)$ in case $|N|=4$ contains 23 1902319023\,19023 190 models, which fall down into $1352$ permutational equivalence classes. The number of (meet) irreducible elements in this lattice $(\mathfrak{sg}^{\diamond}(N),\subseteq)$ is 385 and they break into 29 permutational types. Note that, in comparison with $\mathfrak{sg}(N)$, the lattice is smaller but the number of its irreducible elements is higher.
The lattice $(\mathfrak{sg}^{\diamond}(N),\subseteq)$ has 31 coatoms in this case, which break into 9 types. Therefore, it is not coatomistic. There is only one permutational type of semi-graphoidal coatoms which is not self-adhesive. It leads to an example of a CI model in $\mathfrak{st}(N)\setminus\mathfrak{sg}^{\diamond}(N)$, the largest one in case $|N|=4$.
Example 5.1.
( $\mathfrak{st}(N)\setminus\mathfrak{sg}^{\diamond}(N)\neq\emptyset$ )
Consider $N:=\{a,b,c,d\}$ and put
:=${\mathcal{M}}:=\{\,\,ab|c,\,ab|d,\,ab|cd,\,\,\,ac|b,\,ad|b,\,bc|a,\,bd|a,\,\,\,cd|\emptyset,\,cd|a,\,cd|b\,\,\}\,.$
This model ${\mathcal{M}}$ is a coatom both in $(\mathfrak{sg}(N),\subseteq)$ and $(\mathfrak{st}(N),\subseteq)$ (which elements coincide in case $|N|=4$). Nevertheless, ${\mathcal{M}}\not\in\mathfrak{sg}^{\diamond}(N)$. To observe this we show below that ${\mathcal{M}}$ is not self-adhesive at $L:=bd$ relative to $\mathfrak{sg}$. We apply Lemma 3.4 with $\mathfrak{f}:=\mathfrak{sg}$ for this purpose. Fix a bijection $\varphi:N\to M:=\{\bar{a},b,\bar{c},d\}$, where $\varphi(a):=\bar{a}$, $\varphi(b):=b$, $\varphi(c):=\bar{c}$, and $\varphi(d):=d$. To show that $bc|\emptyset\in\mathfrak{sg}^{\diamond}({\mathcal{M}}|L)\setminus{\mathcal{M}}$ it is enough to verify that $bc|\emptyset\in\mathfrak{sg}^{\diamond}({\mathcal{R}}|L)$ for a subset :=${\mathcal{R}}:=\{\,ab|d,\,ad|b,\,bc|a,\,bd|a,\,cd|\emptyset\,\}$ of ${\mathcal{M}}$.
By Lemma 3.4, this means that $bc|\emptyset$ belongs to the semi-graphoidal closure of the union of ${\mathcal{R}}$, $\varphi({\mathcal{R}})$, and $[ac,\bar{a}\bar{c}|bd]$. The following is one of possible semi-graphoidal derivation sequences:
  1. (1)
    $ad|b\in{\mathcal{R}},[a,\bar{a}|bd]~{}\to~{}ad|\bar{a}b~{}\&~{}bd|\bar{a}\in\varphi({\mathcal{R}})~{}\to~{}ad|\bar{a}$,
  2. (2)
    $ab|d\in{\mathcal{R}},[a,\bar{a}\bar{c}|bd]~{}\to~{}[a,b\bar{c}|\bar{a}d]~{}\&~{}ad|\bar{a}\,\mbox{(1.)}~{}\to~{}ab|\bar{a}\bar{c}$,
  3. (3)
    $\bar{a}d|b\in\varphi({\mathcal{R}}),[\bar{a},ac|bd]~{}\to~{}\bar{a}c|ab~{}\&~{}bc|a\in{\mathcal{R}}~{}\to~{}bc|a\bar{a}$,
  4. (4)
    $\bar{a}b|d\in\varphi({\mathcal{R}}),[\bar{a},c|bd]~{}\to~{}\bar{a}c|d~{}\&~{}cd|\emptyset\in{\mathcal{R}}~{}\to~{}\bar{a}c|\emptyset$,
  5. (5)
    $b\bar{c}|\bar{a}\in\varphi({\mathcal{R}})~{}\&~{}ab|\bar{a}\bar{c}\,\mbox{(2.)}~{}\to~{}ab|\bar{a}~{}\&~{}bc|a\bar{a}\,\mbox{(3.)}~{}\to~{}bc|\bar{a}~{}\&~{}\bar{a}c|\emptyset\,\mbox{(4.)}~{}\to~{}bc|\emptyset$.
Remark that the derivation like $ab|d~{}\&~{}[a,\bar{a}\bar{c}|bd]~{}\to~{}[a,b\bar{c}|\bar{a}d]$ is based on the application of semi-graphoidal inference in its global mode (see Definition 2.9). In this case, this is a shorthand for two-stage derivation:
$[\,a\perp\!\!\!\perp b\,|\,d~{}\&~{}a\perp\!\!\!\perp\bar{a}\bar{c}\,|\,bd\,]~{}\rightarrow~{}a\perp\!\!\!\perp b\bar{a}\bar{c}\,|\,d~{}\rightarrow~{}a\perp\!\!\!\perp b\bar{c}\,|\,\bar{a}d\,.$
Note that the only set $L$ at which the model ${\mathcal{M}}$ is self-adhesive is $cd$.
Of course, the derivation sequence from Example 5.1 also confirms that the considered semi-graphoid ${\mathcal{R}}$ is not self-adhesive at $L=bd$. This is a kind of a certificate (= a direct proof) of the validity of a later implicational rule (SI:3) for self-adhesive semi-graphoids; take $i:=b$, $j:=a$, $k:=d$, and $\ell:=c$ in (SI:3).
The simplest example of a self-adhesive semi-graphoid that is not structural is given in Example 2.8. Its minor modification below shows that $\mathfrak{sg}^{\diamond}$ is not closed under duality, analogously to the $\mathfrak{pr}$-frame (see Example 2.5).
Example 5.2.
( $\mathfrak{sg}^{\diamond}(N)\setminus\mathfrak{st}(N)\neq\emptyset$,  $\mathfrak{sg}^{\diamond}$ is not closed under duality )
Put $N:=\{a,b,c,d\}$ and consider the CI model :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b,\,bc|d\,\}$. It appears to be a self-adhesive semi-graphoid which is not structural: apply the rule (E:1) presented in Section 5.2 with $i:=a$, $j:=b$, $k:=c$, and $\ell:=d$. Nevertheless, its dual model =${\mathcal{M}}^{\star}=\{\,ab|d,\,ac|b,\,ad|c,\,bc|a\,\}$ is not self-adhesive. This follows from the below-mentioned property (SE:1) of self-adhesive semi-graphoids, where $i:=a$, $j:=c$, $k:=b$, and $\ell:=d$. Note that ${\mathcal{M}},{\mathcal{M}}^{\star}\not\in\mathfrak{st}(N)$ and, thus, they are not probabilistic CI models.
Note in this context that, if $|N|=4$ then 48 permutational types of models ${\mathcal{M}}\in\mathfrak{sg}^{\diamond}(N)$ with ${\mathcal{M}}^{\star}\not\in\mathfrak{sg}^{\diamond}(N)$ exist.
We also computed the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$ for ${\mathcal{F}}=\mathfrak{sg}^{\diamond}(N)$ in case $|N|=4$. It decomposes into 9 permutational types: besides two semi-graphoidal types mentioned in Section 5.1 it has 7 additional types and we divided them into two subgroups. The implications in the first subgroup can be interpreted as weaker versions of probabilistic equivalences from [31, § V.A], recalled in Section 5.2, and, therefore, we use the label SE to denote them:
  • (SE:1)
    $\{\,ij|k,\,ik|\ell,\,jk|i,\,i\ell|j\,\}\,\to\,\{\,ij|\ell,\,ik|j,\,i\ell|k\,\}$,
  • (SE:2)
    $\{\,ij|k,\,ij|k\ell,\,ik|j,\,jk|i,\,jk|i\ell,\,i\ell|\emptyset,\,i\ell|j,\,k\ell|\emptyset\,\}\,\to\,\{\,k\ell|j\,\}$.
Indeed, the implication (SE:1) weakens the property (E:1) with an additional antecedent $jk|i$, while the rule (SE:2) weakens the rule (E:5): the 4 additional antecedents are $ij|k\ell$, $ik|j$, $jk|i$, and $i\ell|\emptyset$.
The implications in the second subgroup can be interpreted as weaker versions of probabilistic implications presented in [31, § V.B], for which reason we use the label SI to denote them:
  • (SI:1)
    $\{\,ij|k,ik|j,jk|i,jk|\ell,i\ell|j\,\}\to\{\,ij|\ell\,\}$,
  • (SI:2)
    $\{\,ij|k,ik|j,jk|i,jk|i\ell,i\ell|j\,\}\to\{\,ij|k\ell\,\}$,
  • (SI:3)
    $\{\,ij|k,ik|j,jk|i,i\ell|j,k\ell|\emptyset\,\}\to\{\,i\ell|\emptyset\,\}$,
  • (SI:4)
    $\{\,ij|k,ik|j,jk|i,i\ell|j,k\ell|i\,\}\to\{\,i\ell|k\,\}$,
  • (SI:5)
    $\{\,ij|k,ij|\ell,ik|j,jk|i,k\ell|\emptyset\,\}\to\{\,ij|\emptyset,ik|\emptyset,jk|\emptyset\,\}$.
This is a comparison with the implications from [31, § V.B]:
  • remove the antecedent $jk|i$ in (SI:1) to get (I:10) or $ik|j$ to get (I:11),
  • remove $jk|i$ in (SI:2) to get (I:17) or $ik|j$ to get (I:18),
  • remove $ik|j$ in (SI:3) to get (I:5) or $jk|i$ to get (I:6),
  • remove $ik|j$ in (SI:4) to get (I:8) or $jk|i$ to get (I:12),
  • remove $jk|i$ and $jk|\emptyset$ in (SI:5) to get (I:2) or $ij|k$ and $ij|\emptyset$ to get (I:4).
The result of our analysis was also the observation that the lattice $(\mathfrak{sg}^{\diamond}(N),\subseteq)$ is not implicatively perfect in case $|N|=4$. Indeed, the implication (SE:2) belongs to its canonical implicational basis but it is not perfect: its antecedent set is the union of CI statements from rules (SI:2) and (SI:3).
As illustrated by Example 2.21, the rule (SE:2) can further be simplified:
  • $\overline{\mbox{(SE:2)}}$
    $\{\,ij|k,\,ik|j,\,jk|i,\,jk|i\ell,\,i\ell|j,\,k\ell|\emptyset\,\}\,\to\,\{\,k\ell|j\,\}$.
We also confirmed independently the validity of the rules (SE:1)–(SE:2) and (SI:1)–(SI:5) within the frame $\mathfrak{sg}^{\diamond}$ by finding their certificates in the form of semi-graphoidal derivation sequences, as illustrated in Example 5.1. Specifically, we found out that the antecedent set in (SE:1) is not self-adhesive at sets $ij$ and $jk$, while the antecedent sets in (SE:2) and in the rules labeled by SI are not self-adhesive at the sets $ij$, $ik$, and $jk$.
5.4. Self-adhesive structural semi-graphoids: 4 variables
The family $\mathfrak{st}^{\diamond}(N)$ in case $|N|=4$ contains 20 9682096820\,96820 968 models, falling down into $1224$ permutational equivalence classes. The number of (meet) irreducible elements in the lattice $(\mathfrak{st}^{\diamond}(N),\subseteq)$ is 85 and they break into 13 permutational types, presented in the form of a concise catalogue in Appendix D.
The same phenomenon was observed as in Section 5.3: in comparison with $(\mathfrak{st}(N),\subseteq)$, the number of elements of the lattice decreases while the number of its irreducible elements increases. The set of coatoms of $(\mathfrak{st}^{\diamond}(N),\subseteq)$ is the same as in case of $(\mathfrak{sg}^{\diamond}(N),\subseteq)$, namely 31 of them falling into 9 permutational types. In particular, the lattice $(\mathfrak{st}^{\diamond}(N),\subseteq)$ is not coatomistic for $|N|=4$.
We also computed the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$ for ${\mathcal{F}}=\mathfrak{st}^{\diamond}(N)$ in case $|N|=4$. It decomposes into 12 permutational types: it consists of two semi-graphoidal types mentioned in Section 5.1, five types (E:1)–(E:5) from Section 5.2 and five types (SI:1)–(SI:5) presented in Section 5.3. It makes no problem to observe that all these implications are perfect, implying that the lattice $(\mathfrak{st}^{\diamond}(N),\subseteq)$ is implicatively perfect in case $|N|=4$.
Our computation also revealed 30 types of models ${\mathcal{M}}\in\mathfrak{st}^{\diamond}(N)$ such that ${\mathcal{M}}^{\star}\not\in\mathfrak{st}^{\diamond}(N)$. One of these examples is as follows.
Example 5.3.
( $\mathfrak{st}^{\diamond}$ is not closed under duality )
Put $N:=\{a,b,c,d\}$. The model :=${\mathcal{M}}:=\{\,ab|d,ac|d,ad|b,bc|\emptyset,bc|d\,\}$ belongs to $\mathfrak{st}^{\diamond}(N)$ while :=${\mathcal{M}}^{\star}:=\{\,ab|c,ac|b,ad|c,bc|a,bc|ad\,\}\not\in\mathfrak{st}^{\diamond}(N)$. The latter fact follows from the rule (SI:2) above where $i:=a$, $j:=c$, $k:=b$, and $\ell:=d$.
Figure 4. Relation between considered families of CI models in case $|N|=4$.
A kind of surprise for us was the observation that $\mathfrak{sg}^{\diamond}(N)\cap\mathfrak{st}(N)=\mathfrak{st}^{\diamond}(N)$ if $|N|=4$ (see Figure 4 for illustration); thus, self-adhesivity relative to $\mathfrak{sg}$ and relative to $\mathfrak{st}$ yield the same restriction. Therefore, we have a natural question.
{open}
Find our whether the relation $\mathfrak{sg}^{\diamond}(N)\cap\mathfrak{st}(N)=\mathfrak{st}^{\diamond}(N)$ hold in general.
5.5. The case of 5 variables
We had the following special goal in this case. There are $1319$ permutational types of coatoms of the lattice $(\mathfrak{st}(N),\subseteq)$ of structural semi-graphoids over 5 variables. Note that the overall number of coatoms of this lattice is 117 978117978117\,978117 978.
In general, the coatoms of this lattice correspond to the extreme rays of the cone of standardized supermodular functions, that is, to supermodular functions $m:{\mathcal{P}}(N)\to{\mathbb{R}}$ satisfying $m(K)=0$ if $K\subseteq N$, $|K|\leq 1$, see [33, Corollary 30]. Thus, the coatoms for $|N|=5$ were obtained by computing the extreme rays of that cone [32]. We raised a couple of questions:
  • Which of the permutational types belong to $\mathfrak{sg}^{\diamond}(N)$?
  • Which of them belong to $\mathfrak{st}^{\diamond}(N)$?
Surprising answers were that only 154 of those types belong to $\mathfrak{sg}^{\diamond}(N)$ and these coincide with the types belonging to $\mathfrak{st}^{\diamond}(N)$. This observation seems to support the conjecture indicated in Task 5.4. Consequences of those results are discussed in detail in Section 6.
5.6. Other examples of abstract CI frames
Researchers interested in graphical statistical models introduced various other abstract CI frames [26, 27]; here we follow the line from [18]. A graphoid over a variable set $N$ is a semi-graphoid ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ satisfying the implication
$ij|\ell K,\,i\ell|jK\in{\mathcal{M}}~{}~{}\Rightarrow~{}~{}ij|K,\,i\ell|K\in{\mathcal{M}}\,.$
A semi-graphoid is called compositional if it satisfies the reversed implication, and a compositional graphoid is a semi-graphoid satisfying both implications. It is evident that ${\mathcal{M}}$ is a graphoid if and only if ${\mathcal{M}}^{\star}$ is a compositional semi-graphoid.
The above definitions, thus, delimit the frame $\mathfrak{gr}$ of graphoids, the frame $\mathfrak{gr}^{\star}$ of compositional semi-graphoids, and the frame $\mathfrak{cgr}$ of compositional graphoids. By definition, $\mathfrak{cgr}\subseteq\mathfrak{gr},\mathfrak{gr}^{\star}\subseteq\mathfrak{sg}$. Our notation reflects the fact that $\mathfrak{gr}^{\star}$ can be interpreted as the dual frame to $\mathfrak{gr}$ because ${\mathcal{M}}\in\mathfrak{gr}^{\star}(N)\,\Leftrightarrow\,{\mathcal{M}}^{\star}\in\mathfrak{gr}(N)$ for any variable set $N$ and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$. It is known that $\mathfrak{gr}\setminus\mathfrak{pr}\neq\emptyset$ and $\mathfrak{pr}\setminus\mathfrak{gr}\neq\emptyset$. In case $|N|=4$, precisely $369$ permutational types of graphoids are probabilistically representable.
A result of our computations was the finding that, in case $|N|=4$, the family $\mathfrak{gr}(N)$ contains 6482 elements, which fall down into 421 permutational types. The number of (meet) irreducible models in $\mathfrak{gr}(N)$ is 408 and they decompose into 32 types. The lattice $(\mathfrak{gr}(N),\subseteq)$ has 14 coatoms of 4 types; thus, it is not coatomistic. The family $\mathfrak{gr}^{\diamond}(N)$ in case $|N|=4$ coincides with $\mathfrak{gr}(N)$.
Since, for any $N$, $(\mathfrak{gr}^{\star}(N),\subseteq)$ is order-isomorphic to $(\mathfrak{gr}(N),\subseteq)$, the numbers of elements, irreducible elements, and coatoms of $(\mathfrak{gr}^{\star}(N),\subseteq)$ are the same as in case of graphoids. What is, however, different is that $\mathfrak{gr}^{\star}(N)\not=(\mathfrak{gr}^{\star})^{\diamond}(N)$ in case $|N|=4$. The lattice $(\mathfrak{gr}^{\star})^{\diamond}(N)$ has only 6182 (instead of 6482) elements in 404 permutational types. It has 14 coatoms in 4 types and 428 irreducible elements falling to 33 types. This is the first concrete example of a frame $\mathfrak{f}$ such that both $\mathfrak{f}$ and $\mathfrak{f}^{\star}$ admit self-adhesion and $(\mathfrak{f}^{\star})^{\diamond}\not=(\mathfrak{f}^{\diamond})^{\star}$, i.e., the operations of duality and self-adhesion on a CI frame do not commute in general.
The family $\mathfrak{cgr}(N)=\mathfrak{gr}(N)\cap\mathfrak{gr}^{\star}(N)$ has 2084 elements in case $|N|=4$ which fall into 157 types. It has 470 irreducible models of 30 types and 6 coatoms of 1 type. In particular, the lattice is not coatomistic. Of course, $\mathfrak{cgr}(N)\subseteq\mathfrak{gr}(N)\subseteq\mathfrak{sg}^{\diamond}(N)$ in this case. Like $\mathfrak{gr}(N)$ in case $|N|=4$, $\mathfrak{cgr}(N)=\mathfrak{cgr}^{\diamond}(N)$.
A quite special CI frame is given by undirected graphical models, which were called separation graphoids in [18]. Given an undirected graph $G$ over a variable set $N$ (= a graph having $N$ as the node set), we say that $ij|K\in{\mathcal{K}}(N)$ is represented in $G$ (after the separation criterion) and write $i\perp\!\!\!\perp j\,|\,K\,\,[G]$ if every path in $G$ between $i$ and $j$ contains a node in $K$. Then $G$ induces a CI model
${\mathcal{M}}_{G}~{}:=~{}\{\,ij|K\,:\ i\perp\!\!\!\perp j\,|\,K\,\,[G]\,\}\subseteq{\mathcal{K}}(N)\,.$
The family of undirected graphical models is then
$\mathfrak{ug}(N)~{}:=~{}\{\,{\mathcal{M}}_{G}\,:\ \mbox{$G$ is an undirected graph over $N$}\,\}\,.$
This defines a CI frame $\mathfrak{ug}$, which is closed under copying and marginalization, but not under intersection. Well-known facts are that $\mathfrak{ug}\subseteq\mathfrak{pr}\subseteq\mathfrak{sg}$ [14]. The $\mathfrak{ug}$-frame has a finite “axiomatization” [26]. We can now give an example that illustrates the claim in the remark n. 2 after Definition 3.3, namely that there is a pair of frames $\mathfrak{f}_{1}\subseteq\mathfrak{f}_{2}\subseteq\mathfrak{sg}$ closed under marginalization and copying and a model ${\mathcal{M}}\in\mathfrak{f}_{1}(N)$ which is self-adhesive relative to $\mathfrak{f}_{2}$ but not relative to $\mathfrak{f}_{1}$.
Example 5.4.
( self-adhesivity is relative to a frame )
Consider frames $\mathfrak{f}_{1}:=\mathfrak{ug}$, $\mathfrak{f}_{2}:=\mathfrak{sg}$ and the variable set $N:=\{a,b,c\}$. The model :=${\mathcal{M}}:=\{\,bc|a\,\}\subseteq{\mathcal{K}}(N)$ belongs to $\mathfrak{ug}(N)$: it is induced by the graph $G$ obtained from the complete graph over $N$ by the removal of the edge between $b$ and $c$. Then ${\mathcal{M}}\in\mathfrak{sg}^{\diamond}(N)$ but ${\mathcal{M}}\not\in\mathfrak{ug}^{\diamond}(N)$. To this end consider the set $L:=bc$ and a bijection $\varphi:N\to M:=\{\bar{a},b,c\}$ with $\varphi(a):=\bar{a}$, $\varphi(b):=b$, and $\varphi(c):=c$ (see Figure 2). Then ${\mathcal{M}}$ is not self-adhesive at $L$ relative to $\mathfrak{ug}$. Indeed, a potential adhesion ${\mathcal{E}}$ of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$ in $\mathfrak{ug}(NM)$ has to satisfy
$\{\,bc|a,\,bc|\bar{a},\,a\bar{a}|bc\,\}\subseteq{\mathcal{E}}~{}~{}\mbox{and}~{}~{}\{ab|c,\,ac|b,\,\bar{a}b|c,\,\bar{a}c|b\,\}\cap{\mathcal{E}}=\emptyset\,,$
and there is no undirected graph $H$ over $NM$ with ${\mathcal{M}}_{H}={\mathcal{E}}$.
Another CI frame, which is not discussed in this paper, consists of regular Gaussian CI models [2], that is, CI structures induced by (continuous) Gaussian probability distributions. This frame is closed under copying, marginalization, and duality but not under intersection.
The frame of gaussoids was introduced in [18] and examined in [4]. This frame approximates the frame of (regular) Gaussian CI models. A recent paper [3] deals with the structural self-adhesivity concept for gaussoids and applies these techniques to Gaussian random variables.
6. Application to entropic region delimitation
[…]
Well-known facts are that the entropy function $h_{\mbox{\scriptsize\boldmath$\xi$}}:{\mathcal{P}}(N)\to{\mathbb{R}}$ is submodular, non-decreasing with respect to inclusion, and satisfies $h_{\mbox{\scriptsize\boldmath$\xi$}}(\emptyset)=0$. Thus, in terms of Section 2.2.3, it is (the rank function of) a polymatroid over $N$.
A polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is then called entropic if $h=h_{\mbox{\scriptsize\boldmath$\xi$}}$ for some discrete random vector $\xi$ over $N$. It is called almost entropic if it is a limit of entropic polymatroids (in the Euclidean topology on ${\mathbb{R}}^{{\mathcal{P}}(N)}$). The entropic region, also named the entropy region, is the set of all entropic polymatroids and the almost entropic region is the set of all almost entropic polymatroids; the latter set is known to be a closed cone in ${\mathbb{R}}^{{\mathcal{P}}(N)}$ [35, § 15.1].
Information inequalities, that is, linear inequalities valid for vectors in the (almost) entropic region, are one of the research topics in information theory. In this paper we discuss unconstrained information inequalities, valid for almost entropic vectors. Besides classic Shannon’s inequalities, defining the cone of polymatroids, a number of independent such inequalities has been found; the first of them in 1998 [37]. The techniques to derive these inequalities follow the scheme of the so-called copy lemma [10] and inspired the concept of a self-adhesive polymatroid by Matúš [23], recalled below in analogy with Definition 3.3.
Definition 6.1 (self-adhesive polymatroid).
Given a bijection $\varphi:N\to M$, the $\varphi$-copy of a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is the polymatroid $h\circ\varphi_{-1}\in{\mathbb{R}}^{{\mathcal{P}}(M)}$. A pair of sets $N,M$ is called modular in a polymatroid $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ if $\hat{h}(N)+\hat{h}(M)=\hat{h}(NM)+\hat{h}(N\cap M)$. A polymatroid $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ extends $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ if $h$ is the restriction of $\hat{h}$ to ${\mathcal{P}}(N)$. We say that a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is self-adhesive at a set $L\subseteq N$ if,
$\forall\,\varphi:N\to M~{}\mbox{bijection with $L=N\cap M$ and $\varphi_{\,|L}=\mbox{\rm id}_{L}$~{}}$
[…]
such that N,M is a modular pair in $~{}~{}\mbox{such that $N,M$ is a modular pair in $\hat{h}$}\,.$
A polymatroid over $N$ is self-adhesive if it is self-adhesive at any $L\subseteq N$.
Recall from Section 2.2.3 that a polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ induces a CI model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ as follows: $ij|K\in{\mathcal{M}}$ if and only if $\Delta h(ij|K)=0$.
Lemma 6.2.
The following statements are true.
  • (a)
    The set of self-adhesive polymatroids is a polyhedral cone in ${\mathbb{R}}^{{\mathcal{P}}(N)}$.
  • (b)
    Every almost entropic polymatroid is self-adhesive.
  • (c)
    Every self-adhesive polymatroid $h$ over $N$ induces a CI model in $\mathfrak{st}^{\diamond}(N)$.
Proof 6.3.
To show (a) is true, consider a set $L\subseteq N$ and the respective bijection $\varphi:N\to M$. Realize that the set of polymatroids $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ in which $N,M$ is a modular pair and that satisfy $\hat{h}(A)=\hat{h}(\varphi(A))$ for $A\subseteq N$ is a polyhedral cone. A polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$ is self-adhesive at $L$ if and only if it belongs the projection of that set of vectors $-^\hat{h}$ from ${\mathbb{R}}^{{\mathcal{P}}(NM)}$ to ${\mathbb{R}}^{{\mathcal{P}}(N)}$. Basic facts from polyhedral geometry are that the class of polyhedral cones is closed under projections and finite intersections. The application of these facts yields (a).
For (b) we first realize that any entropic polymatroid is self-adhesive. The arguments for this from [23, Corollary 1] are analogous to those in our proof of Lemma 3.7. Then (a) implies that set of self-adhesive polymatroids over $N$ is closed, which forces that it contains the closure of the entropic region.
As concerns (c), recall from the remark n. 1 in the end of Section 2.2.3 that ${\mathcal{M}}\in\mathfrak{st}(N)$ if and only if it is induced by a polymatroid over $N$. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ induced by a self-adhesive polymatroid $h\in{\mathbb{R}}^{{\mathcal{P}}(N)}$, one needs to observe that ${\mathcal{M}}$ is self-adhesive at (any) $L\subseteq N$ relative to $\mathfrak{st}$. Having fixed the corresponding bijection $\varphi:N\to M$, take the polymatroid $\hat{h}\in{\mathbb{R}}^{{\mathcal{P}}(NM)}$ from Definition 6.1 and it is straightforward to verify that it induces ^$\hat{{\mathcal{M}}}\in\mathfrak{st}(NM)$ which is an adhesive extension of ${\mathcal{M}}$ and $\varphi({\mathcal{M}})$.
Lemma 6.2 has the following consequence.
Corollary 7.
If a polymatroid $h$ over $N$ induces a CI model ${\mathcal{M}}$ outside $\mathfrak{st}^{\diamond}(N)$ then there is a valid (unconstrained) information inequality which is not true for $h$ and none of positive multiples of $h$ is almost entropic.
Proof 6.4.
Realize that any facet-defining inequality for the cone of self-adhesive polymatroids is a valid information inequality.
To interpret the results from Section 5.5 we recall some geometric facts. A polymatroid $h$ over $N$, $|N|\geq 1$, is called tight if $h(N)=h(N\setminus i)$ for any $i\in N$ and modular if $\Delta h(ij|K)=0$ for any $ij|K\in{\mathcal{K}}(N)$. Every polymatroid $h$ then decomposes into its modular part and its tight part $\tilde{h}$ inducing the same CI model as $h$. Moreover, there exists a linear one-to-one mapping between tight polymatroids and standardized supermodular functions which preserves the induced CI model, see [33, formulas (32) in § 7.2]. This allows one to observe, using Corollary 30 in [33, Appendix B], that a polymatroid $h$ over $N$ generates a non-modular (= tight) extreme ray of the polymatroidal cone if and only if it induces a coatom in the lattice $(\mathfrak{st}(N),\subseteq)$.
Thus, in light of Corollary 7, the computational findings from Section 5.5 imply that most of the 117 978117978117\,978117 978 extreme rays of the cone of tight polymatroids over 5-element variable set are outside the almost entropic region.
Being motivated by this perspective, we put (for every variable set $N$):
$\overline{\mathfrak{pr}}(N)~{}:=~{}\{\,{\mathcal{M}}\subseteq{\mathcal{K}}(N)\,:\ \mbox{${\mathcal{M}}$ is induced by}$
$\hskip 102.42992pt\mbox{an almost entropic polymatroid over $N$}\,\}\,,$
which defines the almost entropic CI frame $\overline{\mathfrak{pr}}$.
{open}
Find out whether one has $\mathfrak{pr}=\overline{\mathfrak{pr}}$ or not.
It is easy to see that the frame $\overline{\mathfrak{pr}}\subseteq\mathfrak{sg}$ is closed under marginalization, copying, intersection, and lifting. Since ${\mathcal{K}}(N)\in\mathfrak{pr}(M)\subseteq\overline{\mathfrak{pr}}(M)$ if $N\subseteq M$, one, thus, has $\overline{\mathfrak{pr}}(N)\subseteq\overline{\mathfrak{pr}}(M)$ then. Therefore, in light of Definition 2.25, Task 6 is equivalent to the question whether the probabilistic CI implication on ${\mathcal{K}}(N)$ coincides with the (almost entropic) $\overline{\mathfrak{pr}}$-closure operation (for any $N$). This was basically a question raised in a private communication by our colleague Rostislav Matveev. We incline to believe that there are valid CI implications which are not true in the frame $\overline{\mathfrak{pr}}$.
The suspected probabilistic CI rules from [31, § V.B] which may appear to be examples of invalid implications in the $\overline{\mathfrak{pr}}$-frame are the rules (I:1), (I:7), (I:13), and (I:15). The reason for this guess is that to derive these implications one inevitably needs the first conditional Ingleton inequality [36] and this particular conditional inequality need not hold for almost entropic vectors [17, § V.1].
{remarks}
Let us add two remarks.
  1. (1)
    It follows from Corollary 7 that any CI implication valid in $\mathfrak{st}^{\diamond}$-frame can be viewed as a consequence of an unconstrained information inequality. Note, however, that conditional information inequalities are also studied in information theory, whose validity is restricted to and guaranteed for entropic vectors only [17]. All the so-far known probabilistic CI implications were derived from such conditional inequalities [31].
  2. (2)
    We also asked ourselves whether some of the remaining 154 types of extreme polymatroids can be excluded from being entropic using stronger probabilistic rules from [31, § V.B]. This is indeed the case: we found further 11 types which cannot be probabilistic because their induced CI models ${\mathcal{M}}$ do not satisfy the (extended version of the) probabilistic rule (I:2) saying that
    $ij|M,\,ik|\ell M,\,k\ell|iM,\,k\ell|jM\,\in{\mathcal{M}}~{}\Rightarrow~{}ik|M,\,k\ell|M\,\in{\mathcal{M}}\,.$
    But only 5 of these 11 types can reliably be excluded from being almost entropic as they do not satisfy the above implication with $M=\emptyset$; these finer conclusions follow from the (method of) rule derivation in [31, § V.B].
7. Conclusions
[…]
{open}
Find all coatoms and irreducible elements of $(\mathfrak{sg}(N),\subseteq)$ in case $|N|=5$.
Note in this context that, according to a counter-example from [16], in case $|N|=5$, coatoms of $(\mathfrak{sg}(N),\subseteq)$ exist which are outside $\mathfrak{st}(N)$.
We also verified computationally that $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}^{\diamond\diamond}(N)$ in case $|N|=4$. Computing this “second-order self-adhesion” required a significant implementation effort and took much longer than any of our other computations. Nonetheless, we expect that $\mathfrak{sg}^{\diamond}(N)=\mathfrak{sg}^{\diamond\diamond}(N)$ does not hold for $|N|=5$.
[…]
We end with an open question about finite axiomatizability. We have seen in Section 5.1 that the semi-graphoidal frame $\mathfrak{sg}$ is special in the sense that the canonical implicational basis for $(\mathfrak{sg}(N),\subseteq)$ for any $|N|$ follows one scheme only. One can interpret this by saying that semi-graphoids have a finite “axiomatization” (with a single axiom).
{open}
Find out whether $\mathfrak{sg}^{\diamond}$ admits a finite “axiomatic” characterization, too.
Acknowledgements
Most of this work was inspired by the research of our late colleague Fero Matúš. The authors discussed in person the topic of this paper at two meetings in 2022, one of them was the DAGSTUHL seminar 22301 “Algorithmic Aspects of Information Theory”.
Appendix A Algorithm to compute the canonical basis
This is a pseudo-code of an algorithm to get the canonical generator ${\mathbb{G}}_{{\mathcal{F}}}$ on basis of a Moore family ${\mathcal{F}}$ of subsets of $X$ or its closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$:
  1. (1)
    choose the maximal parameter $m$, $-1\leq m\leq|X|$, such that any set $Y\subseteq X$ with $|Y|\leq m$ is closed (that is, $Y\in{\mathcal{F}}$); put ${\mathbb{G}}:=\emptyset\subseteq{\mathcal{P}}(X)\times{\mathcal{P}}(X)$,
  2. (2)
    unless $m=|X|$ raise $m:=m+1$; if $m=|X|$ then put ${\mathbb{G}}_{{\mathcal{F}}}:={\mathbb{G}}$ and exit,
  3. (3)
    test $Y\subseteq X$ with $|Y|=m$ whether they are pseudo-closed (Definition 2.20); [^10] Note that pseudo-closed strict subsets of $Y$ are encoded in current version of ${\mathbb{G}}$.
    enlarge ${\mathbb{G}}:={\mathbb{G}}\cup\{\,(Y\to\mbox{\rm cl}_{\mathcal{F}}(Y)\setminus Y)\,:\ Y\in\wp({\mathcal{F}}),\,|Y|=m\,\}$,
  4. (4)
    then compute ${\mathbb{G}}^{\triangleright}$ and branch:
    • (a)
      if $|{\mathbb{G}}^{\triangleright}|>|{\mathcal{F}}|$ then go to 2,
    • (b)
      if $|{\mathbb{G}}^{\triangleright}|=|{\mathcal{F}}|$ then ${\mathbb{G}}_{{\mathcal{F}}}:={\mathbb{G}}$ and exit.
Note that in the step 1, when ${\mathcal{F}}=\mathfrak{f}(N)$ for a suitable CI family $\mathfrak{f}$, one typically has $m=1$. By the construction of ${\mathbb{G}}$, we know that ${\mathbb{G}}\subseteq{{\mathcal{F}}}^{\triangleleft}$ and ${{\mathcal{F}}}\subseteq{\mathbb{G}}^{\triangleright}$ in every phase of the algorithm. In the step 4(b), we are sure that ${\mathbb{G}}$ is the implicational generator of ${\mathcal{F}}$ by the criterion mentioned above Example 2.19.
Appendix B Algorithm to test iterated self-adhesivity
Consider a frame $\mathfrak{f}$ closed under copying, marginalization, intersection, lifting and tight replication, such that $\mathfrak{f}\subseteq\mathfrak{sg}$.
The basis of the algorithm is a subroutine to compute the $\mathfrak{f}^{\diamond}$-closure of a given model ^$\hat{{\mathcal{R}}}\in\mathfrak{f}(M)$, which is a modification of the procedure in Lemma 4.6.
Subroutine $\mathfrak{f}^{\diamond}(\hat{{\mathcal{R}}})$ [the input is ^$\hat{{\mathcal{R}}}\in\mathfrak{f}(M)$]
  1. (1)
    put $^{\mathcal{R}}:=\hat{{\mathcal{R}}}$
  2. (2)
    apply to ${\mathcal{R}}$ the following cycle governed by $L\subseteq M$, $|L|=2,\ldots,|M|-2$:
    • (a)
      compute the self-adhesive closure $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)$ as described in Lemma 3.4,
    • (b)
      test whether $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)\setminus{\mathcal{R}}\neq\emptyset$ :
      • if yes then put ${\mathcal{R}}:=\mathfrak{f}^{\diamond}({\mathcal{R}}|L)$, exit the cycle, and go to 2,
      • if not then continue within the cycle with the next set $L$,
  3. (3)
    put $\mathfrak{f}^{\diamond}(\hat{{\mathcal{R}}}):={\mathcal{R}}$ and exit.
Note that in step 2, by Corollary 4 and Lemma 3.18, we need not test sets $L$ of cardinalities $10,1,|M|-$ 1, and $|M|$. As concerns updating of ${\mathcal{R}}$ in 2(b), by Definition 3.3, $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)\in\mathfrak{f}(M)$.
The main procedure is analogous and differs in small details only. The main difference is that Lemma 3.4 is this time applied to $\mathfrak{f}^{\diamond}$ instead of $\mathfrak{f}$. Note that this modification is possible thanks to Lemma 4.1.
Main procedure $\mathfrak{f}^{\diamond\diamond}(\hat{{\mathcal{M}}})$ [the input is ^$\hat{{\mathcal{M}}}\in\mathfrak{f}^{\diamond}(N)$, where $|N|=4$]
  1. (1)
    put $^{\mathcal{M}}:=\hat{{\mathcal{M}}}$
  2. (2)
    apply to ${\mathcal{M}}$ the following cycle governed by $E\subseteq N$, $|E|=2,3$:
    • (a)
      compute the self-adhesive closure $\mathfrak{f}^{\diamond\diamond}({\mathcal{M}}|E)$ using Lemma 3.4, where the subroutine is called to compute the $\mathfrak{f}^{\diamond}$-closures,
    • (b)
      test whether $\mathfrak{f}^{\diamond\diamond}({\mathcal{M}}|E)\setminus{\mathcal{M}}\neq\emptyset$ :
      • if yes then put ${\mathcal{M}}:=\mathfrak{f}^{\diamond\diamond}({\mathcal{M}}|E)$, exit the cycle, and go to 2,
      • if not then continue within the cycle with the next set $E$,
  3. (3)
    if $^{\mathcal{M}}\neq\hat{{\mathcal{M}}}$ then ^$\hat{{\mathcal{M}}}\not\in\mathfrak{f}^{\diamond\diamond}(N)$ and its closure is $\mathfrak{f}^{\diamond\diamond}(\hat{{\mathcal{M}}}):={\mathcal{M}}$; exit.
Note that, in step 2, by Lemma 4.1 and Corollary 4, we need not test sets $E$ of cardinalities $0,1$, and $4$. On the other hand, since we are not sure whether $\mathfrak{f}^{\diamond}$ is closed under tight replication (see Task 4) we have to test sets $E$ of cardinality $3$. So, the number of sets $E$ in line 2 is at most 10.
To check double self-adhesivity over $4$ variables one needs to apply the subroutine for $|M|=5,6$, in which case the number of sets $L$ in step 2 of the subroutine is at most 50. Thus, to compute $\mathfrak{f}^{\diamond}({\mathcal{R}}|L)$ one applies the $\mathfrak{f}$-closures over at most 10 variables.
Appendix C Table of our results in case of 4 variables
family models in types irreducibles in types coatoms in types
$\mathfrak{sg}(N)$ 26 4242642426\,42426 424 $1512$ $181$ $20$ $37$ $10$
$\mathfrak{sg}^{\diamond}(N)$ 23 1902319023\,19023 190 $1352$ $385$ $29$ $31$ $9$
$\mathfrak{st}(N)$ 22 1082210822\,10822 108 $1285$ $37$ $10$ $37$ $10$
$\mathfrak{st}^{\diamond}(N)$ 20 9682096820\,96820 968 $1224$ $85$ $13$ $31$ $9$
$\mathfrak{gr}(N)$ $6482$ $421$ $408$ $32$ $14$ $4$
$\mathfrak{gr}^{\diamond}(N)$ $6482$ $421$ $408$ $32$ $14$ $4$
$\mathfrak{gr}^{\star}(N)$ $6482$ $421$ $408$ $32$ $14$ $4$
$(\mathfrak{gr}^{\star})^{\diamond}(N)$ $6182$ $404$ $428$ $33$ $14$ $4$
$\mathfrak{cgr}(N)$ $2084$ $157$ $470$ $30$ $6$ $1$
$\mathfrak{cgr}^{\diamond}(N)$ $2084$ $157$ $470$ $30$ $6$ $1$
Appendix D Irreducible self-adhesive structural models
Here we give the list of permutational types of (meet) irreducible elements of the lattice $(\mathfrak{st}^{\diamond}(N),\subseteq)$ in case $|N|=4$: [^11] Recall that we apply the notational convention from Definition 2.9.
  • I.
    $\{\,\,[ijk,\ell|\emptyset],\,[ij,k|\emptyset],\,[ij,k|\ell]\,\,\}$,
  • II.
    $\{\,\,[ijk,\ell|\emptyset],\,ij|k,ij|k\ell,\,ik|j,ik|j\ell,\,jk|i,jk|i\ell\,\,\}$,
  • III.
    $\{\,\,[ij,k|\ell],\,[ij,\ell|k],\,ij|k,ij|\ell,ij|k\ell,\,ik|j,\,i\ell|j,\,jk|i,\,j\ell|i,\,k\ell|i,k\ell|j,k\ell|ij\,\,\}$,
  • IV.
    $\{\,\,[ijk,\ell|\emptyset],\,ij|\emptyset,ij|\ell,\,ik|\emptyset,ik|\ell,\,jk|\emptyset,jk|i\ell\,\,\}$,
  • V.
    $\{\,\,[ij,k|\emptyset],\,[ij,\ell|\emptyset],\,ij|\emptyset,ij|k,ij|\ell,\,ik|\ell,\,i\ell|k,\,jk|\ell,\,j\ell|k,\,k\ell|\emptyset,k\ell|i,k\ell|j\,\,\}$,
  • VI.
    $\{\,\,[ij,k|\ell],\,[ij,\ell|k],\,ij|\emptyset,\,ik|\emptyset,\,i\ell|\emptyset,\,jk|\emptyset,\,j\ell|\emptyset,\,k\ell|ij\,\,\}$,
  • VII.
    $\{\,\,ij|\emptyset,ij|k\ell,\,ik|\emptyset,ik|j\ell,\,i\ell|\emptyset,i\ell|jk,\,jk|\emptyset,jk|i\ell,\,j\ell|\emptyset,j\ell|ik,\,k\ell|\emptyset,k\ell|ij\,\,\}$,
  • VIII.
    $\{\,\,[ij,k|\ell],\,ij|\emptyset,ij|\ell,ij|k\ell,\,ik|\emptyset,\,i\ell|jk,\,jk|\emptyset,\,j\ell|ik,\,k\ell|ij\,\,\}$,
  • IX.
    $\{\,\,[ij,k|\emptyset],\,ij|\emptyset,ij|k,ij|k\ell,\,ik|j\ell,\,i\ell|\emptyset,\,jk|i\ell,\,j\ell|\emptyset,\,k\ell|\emptyset\,\,\}$,
  • X.
    $\{\,\,ij|\emptyset,ij|k,ij|\ell,\,ik|\ell,\,jk|\ell,\,k\ell|i,k\ell|j,k\ell|ij\,\,\}$,
  • XI.
    $\{\,\,ij|\emptyset,ij|k,ij|\ell,\,ik|\ell,\,j\ell|k,\,k\ell|i,k\ell|j,k\ell|ij\,\,\}$,
  • XII.
    $\{\,\,ij|\emptyset,ij|k,ij|\ell,\,ik|\ell,\,i\ell|k,\,jk|\ell,k\ell|j,k\ell|ij\,\,\}$,
  • XIII.
    $\{\,\,ij|\emptyset,ij|k,ij|\ell,\,ik|\ell,\,i\ell|k,\,jk|\ell,\,j\ell|k,\,k\ell|ij\,\,\}$.
The types I.IX. involve all 31 coatoms of the lattice. Their notation follows the classification used in [31, Appendix B]. The actual numbers of involved CI statements are as follows:
  • I. consists of 20 CI statements,
  • II.V. consists of 18 CI statements,
  • VI. consists of 14 CI statements,
  • VII.IX. consists of 12 CI statements, and
  • X.XIII. consists of 8 CI statements.